### v1

# Pressure and strain effects on the ab initio electronic structure of La3Ni2O7

Abstract
The recent discovery of superconductivity in La3Ni2O7 at a critical temperature above 80Β K points to a non-conventional pairing mechanism in nickelates as in cuprates, possibly due to electronic correlations. We have calculated from first principles the electronic structure of La3Ni2O7 under the effect of pressure and epitaxial strain including correlations by the approximation to the many-body self-energy. We find that the Fermi surface is composed of a characteristic cuprate-shape sheet plus a nickelate-specific cylinder , both from Ni orbitals, with a non-negligible drop in the quasiparticle weight and an effective 1D character. This topology results from a delicate balance between the Ni-3 hole pocket , which is suppressed by correlations, and an emerging La-5 electron pocket induced by both correlation and pressure/strain effects and whose role at low energy has been neglected so far. Unlike cuprates, the electronic structure of La3Ni2O7 is already correctly described from ab initio and in agreement with the experiment without the need to introduce Hubbard adjustable parameters or to invoke a strongly correlated physics.
I Introduction
The recent discovery of superconductivity (SC) in Ruddlesden-Popper (RP) nickelates [1, 2, 3] has introduced a new direction in the field of high-temperature superconductivity. Notably, superconductivity has been observed in bulk samples of the bilayer La3Ni2O7 under high pressure [1], with the superconducting phase forming a right-triangular region spanning 14 to 80 GPa, and reaching a maximum onset critical temperature of approximately 80Β K at 18 GPa [4]. More recently, superconductivity has also been reported at Β K in La3Ni2O7 thin films under epitaxial compressive strain at ambient pressure [5, 6]. While the superconducting mechanism is still debated in infinite-layer LnNiO2 nickelates [7, 8, 9, 10, 11], where critical temperatures do not overstep 30Β K [12, 13], the high critical temperature observed in La3Ni2O7 is difficult to reconcile with a BCS electron-phonon pairing mechanism [14, 15], thus pointing to unconventional superconductivity like in cuprates. Indeed, cuprates and nickelates share many analogies: they are both organized in a layered crystal structure with CuO2 or NiO2 planes; and present the same βcuprate-shapeβ Fermi surface sheet of main 3 character, although nickelates bilayer and trilayer show a strong Ni-3 hybridization in the antinodal direction. However, there are also important differences: the constant presence of an anti-ferromagnetic insulating parent phase in cuprates has no analogous in nickelates. For what concerns the electronic structure, the fact that in nickelates the Ni-3 states are closer to the Fermi level than in cuprates, gives rise to at least another Fermi sheet and possibly to a multiband character superconductivity. Therefore, although both unconventional, the pairing mechanism could be possibly different in cuprates and nickelates.
Interestingly, unconventional superconductivity was anticipated on a theoretical basis in La3Ni2O7 by invoking finite-energy spin fluctuations as pairing glue [16]. Then many other different model theoretical pictures [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] have been proposed to explain the superconductivity in the bilayer. The symmetry of the superconducting state remains unclear as subsequent studies have indicated the possibility of not only -wave, but also -wave pairing states including -wave, depending on specific details of the electronic structure [24, 34, 49, 50, 51, 31, 52]. Also the respective role of the two Ni-3 orbitals in the superconducting mechanism is still under discussion [53]. With respect to spin fluctuations, the gap between the two upper and lower Ni-3 bands, originating from the interlayer hopping, seems to be relevant [24]. The presence or absence of the so-called hole pocket in the Fermi surface, that should arise from the lower Ni-3 band, is still unclear both in theoretical calculations [45, 54, 55] and in experimental angle-resolved photoemission spectroscopy (ARPES) measurements [56, 57]. Regardless of its presence or not at the Fermi level, this band is believed to be important in the pairing mechanism [58]. All the previous points require a correct description of the electronic structure including electronic correlations.
In this work we study from ab initio the electronic structure of La3Ni2O7 including correlations at the level of the approximation to the self-energy [59, 60, 61, 62]. In contrast to the Kohn-Sham eigenvalues used in density-functional theory (DFT), which are the energies of a fictitious non-interacting system, quasiparticle energies have a direct physical interpretation as the excitation energies measured in ARPES. The approximation takes into account electronic exchange and correlations without introducing any adjustable parameter, neither in the DFT functional (e.g.Β an hybrid mixing parameter) nor in the Hamiltonian (e.g.Β a Hubbard interaction term). The weight of correlation effects and their importance in this system is revealed by a comparison between the electronic structure and the DFT band plot.
We also investigate the electronic structure evolution of bulk La3Ni2O7 with respect to pressure, as well as in presence of in-plane constraints that simulate a thin film grown on a substrate. Our purpose is to highlight aspects which might be relevant to explain superconductivity that in this system is induced by pressure without doping. The focus is at the Fermi level but also at higher energy features which are less affected by the estimated 0.1Β eV best absolute accuracy, and which might be relevant in the competition between different pairing symmetries according to some proposed superconducting mechanisms [24]. Our work describes some of the electronic structure features expected to be relevant by theoretical models so far proposed [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and provides also indications for future works.
The paper is organized as follows: after introducing our methods in Sec.Β II, we will first compare our calculations with previous ab initio works done at 29.5Β GPa with the crystal structure used by almost all DFT as well as two calculations [45, 15] (Sec.Β III.1). We will then show the evolution with pressure of the electronic structure for the bulk crystal in Sec.Β III.2, and finally introduce in Sec.Β III.3 our results for the structure with epitaxial strain of 1.8%. Our results will be compared with the available experimental data from photoemission (ARPES) experiments.
II Calculation details
II.1 Numerical methods
Our methodology relies as a first step on density-functional theory (DFT) calculations in the PBE approximation [63], using the plane waves code Abinit [64]. We use norm-conserving pseudo-potentials from the PseudoDojo table library [65]. The La pseudo-potential takes 4 electrons in conduction and does not freeze them in the core. We have used a 161616 standard sampling of the Brillouin zone (BZ) and a plane waves (PW) cutoff of 65Β Ha for the DFT-PBE starting calculation, as well as a Gaussian smearing of 0.01 Hartree. In the subsequent calculations, the BZ sampling has been reduced to 666, whereas the PW cutoffs have been reduced to 40 Ha for the representation of the wave functions and for the exchange self-energy, and further down to 15 Ha for the correlation self-energy and the screening. We included 187 bands for the calculation of the screening and 225 for the self-energy (see Sect.Β F in the AppendixΒ for convergence studies). We have carried just only one iteration (i.e.Β ) on top of PBE, using a Godby-Needs plasmon-pole model [66] and a shift of 0.1Β eV to avoid poles/divergences. The final band plots, total and projected density of states (DOS and PDOS), as well as Fermi surfaces, were interpolated from the 666 to a denser -mesh by the Wannier90 code using a set of 67 projected Wannier functions (PWFs) with the closest possible character to atomic orbitals, namely O-, Ni-, La- and La-. The band orbital characters and PDOS identified by these PWFs are almost coincident with what one would obtain by projecting Bloch wave functions directly onto real atomic orbitals, which is not the case when selecting a set of maximally-localized Wannier functions (MLWFs) (see App.Β D). In the bandplots showing projected orbital character, the area of the dots scales as the orbital contribution squared, so that the width of the line is proportional to it.
All calculations have been done with a Gaussian smearing temperature of 0.01Β Ha to ease self-consistency convergence and optical absorption onset stability (which in any case has little relevance on main plasmons, energy-loss and screening ). For the most critical case at Β GPa we have checked that a calculation with enforced at Β Ha in both DFT and calculations provides the same physical picture, and in particular the same Fermi surface as in Fig.Β 1b, with only minor adjustments of less than Β meV on band positions, mostly due to the fact that the chemical potential at larger temperature is a less accurate estimate for the Fermi energy (see App.Β C).
II.2 Crystal structure and Brillouin zone
To determine the crystal structure of La3Ni2O7 at ambient pressure, and then in the superconducting range at 14, 29.5 and 40 GPa, we have performed DFT PBE structural relaxations which provided the I4/mmm space group as the lowest energy structure at all considered pressures, except at 0Β GPa where the Amam structure is more stable. Nevertheless, our calculations have all referred to the I4/mmm structure even at ambient pressure to ease the comparison and better follow the evolution of the electronic structure with pressure, from the superconducting range down to 0Β GPa. For all the cases we have used the relaxed lattice parameters and the internal atomic position out of our PBE relaxations, except for the case at 29.5Β GPa which we treated separately.
Indeed, 29.5Β GPa is the most studied case in the theoretical literature, so that we use it as a benchmark case to compare with previous calculations. Thus, for the sake of comparison, we decided to stick to the lattice parameters and atomic positions used in most of previous works [45, 24]. Therefore, we consider a tetragonal structure with Β Γ Β and Β Γ Β as Sakakibara et al.Β [24]. In fact, instead of the initial space-group symmetry proposed for La3Ni2O7 at 29.5Β GPa, our calculations predict the space-group symmetry with lattice parameters that are in good agreement with the experimental ones [67]. Further, in Ref.Β 45, it was noticed that the electronic structure of La3Ni2O7 at 29.5Β GPa is extremely sensitive to internal atomic positions and that the experimental ones reported by Sun et al.Β from the X-ray diffraction [1] likely yield artifacts at the Fermi level. We provide a detailed analysis of this question in AppendixΒ B.
To simulate epitaxial strain, we enforced the tetragonal lattice parameter to the experimentally observed value of 3.77Β Γ Β [5] and let the system relax in the direction, which provided Β Γ . This crystal structure corresponds to a bulk non-hydrostatic planar stressed system, or equivalently a system pulled along the axis, that we will use to simulate an epitaxially strained thin film. All our crystal structures are detailed in App.Β A.
All our electronic structures are reported in the body-centered tetragonal (BCT) Brillouin zone (BZ) which we report in Fig.Β 1c by black lines, while the blue line shows the path followed in our band plots with -point labels using the convention of Hinuma et al. Β [68]. Notice that in this convention, the X and M -points do not correspond to the canonical ones used in the 2D BZ of the square lattice widely used in the model literature: X is at the BZ corner in the nodal direction and M lies in the plane. We also indicate in grey the BZ corresponding to the conventional cell, namely simple tetragonal, which ease the comparison with 2D BZs. In the same Fig.Β 1c, we report also the simple orthorhombic BZ and associated -path to compare with the ARPES experiment [56] which referred to the tilted Amam structure needing a supercell compared to ours; and finally the face-centered orthorhombic BZ (and associated -path) to compare with the calculation of Christiansson et al.Β [45] who referred to the crystal structure.
III Results
III.1 Electronic structure at 29.5 GPa
We start discussing the electronic structure of La3Ni2O7 at the pressure of 29.5Β GPa over which most of previous theoretical calculations have been performed. Our PBE and Fermi surfaces are shown in Fig.Β 1a and 1b, while Fig.Β 1d shows the corresponding band structures. In the same figure, selected orbital character contributions are plotted on the band structure, as well as the projected densities of states (PDOS) and the total density of states (DOS).
We briefly mention that, like it was already found [69, 7, 8] in infinite-layer nickelates, e.g.Β LaNiO2, correlations shift 2Β eV up the flat bands with predominant La-4 character, moving them further away from the Fermi level and out of the low-energy region, and shift 1Β eV down the O-2 manifold, leading to an increased ionic charge-transfer energy between Ni-3 and O-2 states. These aspects were previously detailed [69] and will not be again discussed here where we focus on the low-energy region.
III.1.1 Ni Fermi surface sheets
The Fermi level is dominated by bands with Ni-3 character. The DFT PBE Fermi surface consists of three sheets, labeled , , and in previous literature [18]. The sheet is a large cylindrical sheet centered around ; is the sheet with the characteristic shape of the cuprate Fermi surface; and is a small hole pocket located at the Brillouin zone (BZ) corners which is highly dispersive along . The and sheets are primarily of character, though they show significant hybridization with orbitals in the anti-nodal direction, as well as with O-2 states (see Fig.Β 2b). In contrast, the pocket has an almost pure character and is consistently present in PBE (or LDA) calculations.
The Fermi surface, in contrast, has no hole-pockets, as shown in Fig.Β 1b. This result is totally consistent with experimental ARPES data [56], with respect to which we provide a detailed comparison in Sec.Β III.2 below, since it was performed at ambient pressure. This difference between DFT and is due to a slight downward shift (7080 meV) of the lower Ni-3 band along the XP direction (see Fig.Β 1d). The absence of the pocket in experiment has been discussed also in the literature, and several semi-empirical corrections have been shown to reconcile theoretical calculations with the ARPES experiment. These methods either include the use of a weighted Fock exchange operator, as by HSE [54] or any other hybrid approach; or place more strength on correlation by introducing a Hubbard repulsion term of Β eV on Ni-3 electrons, like it was done in the calculation of Yang et al.Β [56], with the effect to lower this band. On the basis of this fact, Yang et al.Β [56] concluded that La3Ni2O7 is a strongly correlated system. Here we show that, a calculation using physical quasiparticle energies out of a many-body self-energy, even in a non self-consistent approximation, is enough to get rid of the sheet at the Fermi level and achieve good agreement with the experiment. Our results are in line with the findings of Ryee et al.Β [55] who excluded both a transition to a Mott insulator or bad metal scenario.
The disappearance of the hole pocket has also been reported in a previous work [45] (see Fig.Β 1c for -point correspondence). They performed a calculation based on a reduced four-orbital tight-binding model restricted to the Ni- manifold. This model does not include the La-5 band, which, as we will discuss in the next section, plays a non-negligible role in the low energy spectrum. Their model also employs an effective screening that is renormalized to account for the missing degrees of freedom, and refers to the Fmmm structural phase.
Despite these significant methodological differences, our results are in remarkable agreement with theirs regarding the absence of the pocket. However, Christiansson et al.Β do not further comment on this disappearance in their calculation, because their subsequent EDMFT simulationsβperformed on top of both DFT and βreintroduce and even enhance the spectral weight of the lower Ni-3 band at the Fermi level. On this basis, they conclude that βthe dominance of the orbital distinguishes La3Ni2O7 from the infinite-layer nickelates or the cuprates,β which are instead dominated by character. This interpretation is not supported by our results (see Sec.Β III.1.4 for the spectral weights), which show a predominantly character at low energy. Our findings rather align with the recent calculations of You et al.Β [15].
Interestingly, Ryee et al.Β [55] found the pocket in both DFT and dynamical mean-field theory (DMFT) calculations, but not in cluster-DMFT (CDMFT) which is therefore more in agreement with the finding. This is an indication that non-local correlations, as provided by a CDMFT or a -dependent self-energy, are important to describe such modification of the Fermi surface topology. However, CDMFT also find an important modification to the cuprate-shape sheet which loses most of its 1D character (see below). On this point is more in agreement with DFT and local DMFT rather than with CDMFT.
III.1.2 La-5 band lowering and self-doping
Beside this important qualitative modification of the Fermi surface, correlation corrections induce an even more evident and quantitative change in the bandwidth of the 5 band contributed by the two La spacer atoms only (i.e., with no contribution by the La inner atom lying in between the NiO planes, see Fig.Β 13). This La-5 band, which in DFT-PBE is well above the Ni-3 one, is lowered by corrections, as it can be seen at in Fig.Β 1d. At the M point, which is often overlooked in 2D calculations restricted to the plane only, the lowering in is so large (0.7Β eV) that the La-5 crosses the Ni-3 band and becomes the lowest lying conduction band at M. Due to this large correction, the bottom of conduction (BOC) of the La-5 band at M is situated just only 30 meV above the top of valence (TOV) of the Ni-3 band at P (in the following, we will refer to this TOV-BOC distance of 30 meV as the indirect gap). As a consequence, the Fermi level results from a delicate equilibrium between these two incipient bands: the Ni-3 band, which tends to open a pocket of holes at the BZ corners; and the La-5 one, which tends to open a pocket of electrons at M, hereafter labeled . In the latter case, states located on the La atoms of the spacer start to be occupied instead of Ni states. This results in an effective self-doping (SD) by holes of the NiO2 planes. Note that a possible role of rare-earth self-doping was suggested [45] as responsible for the suppression of charge order, thereby possibly favoring superconductivity.
We can expect that any tiny perturbation of the system, e.g.Β doping, vacancies or defects, can drive the system to a qualitatively different electronic structure and Fermi surface, and therefore affect the superconductivity. Note that even computationally, these low-energy features are highly sensitive to the uncertainty of the integration method to calculate the Fermi level, which we discuss in further detail in App.Β C. We mention two recent works which pointed out the shift of the La-5 band due to correlations [15] and highlight the importance of this band on the screening which affects correlations on the Ni- bands [70]. The latter work remarked in particular some coincidences between the lowering trend of the La-5 band with pressure and the shape of the superconducting region. In our calculation this band reaches the Fermi level at lower pressures and crosses the Fermi level already at 30 GPa at the M point, in comparison to 80 Gpa at the point in Verraes et al.Β [70].
III.1.3 Upper-lower Ni-3 energy gap
A feature of the electronic structure that is considered of particular interest in relation to the emergence of superconductivity in La3Ni2O7 is the energy difference between the upper and lower Ni-3 bands. In the literature they are often labeled as bonding- and anti-bonding (as we explain in App.Β H, we find that this labeling is not appropriate). This energy difference is related to the Ni interlayer hopping. According to the scenario proposed by Sakakibara et al.Β [24], a superconductivity gap would open on these two Ni-3 bands even though they can be as far as 1Β eV away from the Fermi level, while the SC gap would be almost zero for the more Ni-3 character bands crossing the Fermi level. In this scenario, a key parameter is precisely this Ni-3 interlayer hopping. As it can be seen in Fig.Β 1d, the two Ni-3 are practically unaffected by corrections, as their band width and separation remain essentially the same as in the DFT calculation. On this point our calculation is in disagreement both with the calculation of Christiansson et al.Β [45], which found a negative downward shift of the upper band of Β eV at the BZ corner, and with the calculation of You et al.Β [15], which in contrast found a positive upward shift of Β eV. Accordingly, our results leave practically unaffected the scenario proposed by Sakakibara et al.Β [24], while it would be boosted or disrupted in the two other cases.
III.1.4 quasiparticle spectral weight and 1D character
A calculation can also provide the quasiparticle spectral weight which represents a measure of the degree of correlation of the system. In Fig.Β 2a we show the interpolated bandplot of La3Ni2O7 at 29.5Β GPa, with the dots corresponding to the energies sampled on the -mesh. The relative size of the dots indicates the corresponding value of . In Fig.Β 2b we show the general orbital character of the bands (for more detailed projections, see Fig.Β 12). These plots are complemented with Fig.Β 3 which illustrates the behavior of as a function of the energy for the N -point of the Brillouin zone, a point we have chosen for the clear picture it gives. We note that La-5 and 4 bands, which at the level are difficult to distinguish, present the maximum QP weight culminating at . This is already quite remarkable, in particular for the most localized flat bands. Some of the La bands, in particular the previously discussed La-5 which almost achieved the Fermi level, bring these high spectral weights lower in energy than the main La manifold located around 5Β eV. For La-5 states above 5 eV, the QP weight is slightly decreasing (see Fig.Β 3). On the other side of the Fermi level, we also observe for the O-2 manifold below Β eV a reduced with increasing distance from with values ranging from 0.75 to 0.70. This is surprisingly lower than for La-5 and 4 states. However, the most remarkable fact is that the Ni-3 bands, with no difference between and , present the lowest weight of all quasiparticle states. For these states we observe a non-negligible drop of 0.15 when compared in particular with the La-5 which at M is situated at the same Fermi level as the Ni-3. This fact points to a larger weight of the non-coherent part of the spectrum for the Ni-3 states.
We remind that in a Fermi liquid [59] the general trend is a reduction of the quasiparticle weight when going from the Fermi level to lower energies. Indeed, when going to the lowest energy quasiparticle states, the non-coherent part of spectra due to, e.g.Β plasmon losses, should increase, and correspondingly decreases. The same happens in the other higher energies direction, although not monotonically at the beginning. From this, we can conclude that already with the correlations introduced by a non-self-consistent approximation, La3Ni2O7 show a Fermi liquid trend for La-5, La-4 and O-2 quasiparticle states, but not for the Ni-3 ones around the Fermi level. The behaviour of the latter shows a departure from the Fermi liquid trend, but the minimal value of is still large enough not to configure a full breakdown of the Fermi liquid towards a strongly-correlated picture.
As shown in Figs.Β 1a and b, the Fermi surface of La3Ni2O7 exhibits negligible dispersion along the direction, highlighting the pronounced 2D character of its electronic structure. We note that, while this situation is similar to what is found in most of the cuprates, it differs from the infinite-layer nickelates case that presents large dispersion which introduces 3D features in their Fermi surface [71, 72]. Moreover, we note that the dispersion of the Fermi surface of La3Ni2O7 along and is also very weak. In fact, the and Fermi surface sheets can effectively be seen as a superposition of two sets of parallel planes along and , respectively. That is, as the superposition of 1D Fermi surfaces. These effectively 1D Fermi surfaces is an important difference with respect to cuprates. In particular, the emergence of or spin-density-wave ordersβas opposed to the antiferromagnetic order characteristic of the cupratesβcan be related to their nesting properties [34].
In the cuprates, ARPES experiments have revealed the emergence of so-called βFermi arcsβ. Specifically, the quasiparticle spectral weight is found to be locally suppressed at points along the Fermi surface that would coincide with the 1D Fermi surface segments. Unfortunately, to check whether a similar phenomenon takes place in La3Ni2O7, the 666 -sampling we have used in our calculation is not precise enough. Nevertheless, any experimental indication along this direction might provide a clue to understand superconductivity in both nickelates and cuprates.
III.2 Evolution with pressure
In order to gain further insight on the key aspects that might be relevant for superconductivity in La3Ni2O7, we also analyze the evolution of its electronic structure with pressure. We remind that La3Ni2O7 enters into a superconducting phase at an experimentally measured pressure of 14Β GPa without doping [1].
To explore the largest range, we have included in our study a fictitious zero pressure phase of La3Ni2O7 enforced in the tetragonal I4/mmm crystal structure, which in reality at 0Β GPa should be unstable towards the orthorhombic Amam phase with tilted octahedra. Of course, at the present knowledge we cannot exclude that the structural transition occurs at the same critical pressure as the transition to superconductivity, so that it is an important ingredient to explain the superconductivity in this system. In this case, our calculation of the tetragonal phase at 0Β GPa would be probably inadequate to study the normal-superconducting phase transition. But beyond the fact that a calculation is less cumbersome in systems with more symmetries, our choice is motivated by the fact that the comparison of electronic structures is difficult between different crystal structures. Topological variations in the Fermi surface might trigger the superconducting phase transition (and maybe also the structural one): the approach proposed in this section focus on them.
In Fig.Β 4 we compare the electronic structure of La3Ni2O7 at ambient pressure and at 29.5Β GPa (in the same figure we also report the band plot of a biaxial strained La3Ni2O7 which will be discussed in the next section). The most remarkable difference between the two pressures is in the La-5 band at the point M which, at ambient pressure, lies eV above the upper Ni-3 band in a very similar way to what already seen but at the DFT-PBE level in the 29.5 GPa case (see Fig.Β 1d), though in the latter the effect is much larger (0.5 eV). That is, correlations and pressure act in the same direction towards the emergence of the La-5 band at the lowest energies. For a more complete discussion on this point, see App.Β B. At ambient pressure the La-5 band is not present in the low energy region, and therefore the delicate balance between the two incipient self-doping bands is lost.
The electronic structure of La3Ni2O7 at ambient pressure has been measured by an ARPES experiment [56] which we can now compare with our calculation at zero GPa. The hole pocket, which is found in the DFT PBE Fermi surface as due to the lower Ni-3 band, is absent in the ARPES experiment, exactly like in (Fig.Β 1b and Fig.Β 16 in the Appendix). ARPES spectra situate the top of this band at the BZ corners and 50Β meV below the Fermi level. Our calculation at zero GPa (see Fig.Β 4), found the top of this band 50Β meV below the Fermi level at the X -point (corresponding to in Yang et al.Β [56], see Fig.Β 1c for -points correspondence). In contrast to the La-5 band, this lower Ni-3 band is almost unaffected by hydrostatic pressure: in the results at 29.5Β GPa it just only shifts down Β meV, which validates the comparison of the latter case with the ARPES, even if not performed at the same pressure.
In Fig.Β 4b, we report a graph presenting the top-of-valence (TOV) and bottom-of-conduction (BOC) positions of the most relevant bands for different hydrostatic pressures. We report extra calculations done at 14 and at 40Β GPa, covering the range of superconductivity as measured experimentally in Sun et al.Β [1]. The indirect gap is at its minimum around 29.5Β GPa, and the La-5 band is clearly self-doping beyond. From the same figure we can also remark that the same La-5 band crosses the Ni-3 upper band precisely at 14Β GPa, which is precisely the onset of the superconducting region measured by Sun et al.Β [1]. This might be just only a coincidence or have deeper implications. In support to the latter, we cite a very recent work [70] which also found a correlation between the position of the La-5 band with respect to the Fermi level, and the shape of the superconducting region. Their calculation is at the DFT level, so that they missed the earlier coincidence in pressure of the La-5 with the Ni-3 band precisely at the onset of superconductivity at 14Β GPa, but the relevance of the La-5 for superconductivity has been remarked. We can more carefully conclude that both correlations and pressure are fundamental to describe the electronic structure of La3Ni2O7 in its superconducting phase. In the next section we will see that -plane strain plays a similar role to correlations and pressure.
III.3 La3Ni2O7 under epitaxial strain
Recent works [5, 6] have reported superconductivity at Β K in La3Ni2O7 thin films under compressive strain at ambient pressure. Here we present the electronic structure of a bulk La3Ni2O7 with a 1.8% strained Β Γ Β and a relaxed , which we use to simulate the experimental thin film. The corresponding PBE and band plots, PDOS and DOS are shown in Fig.Β 5. In comparison to the previously discussed case at ambient and 29.5Β GPa (see also Fig.Β 4), the electronic structure retains most of its features. The Fermi level crossing is practically the same in the three different cases. In particular, the cuprate-like sheet has exactly the same shape (refer to Fig.Β 1). The lower Ni-3 band is significantly further from in the thin film case ( eV). However, we observe two important differences. First, the hole pocket is already absent at the DFT level, that is the lower Ni-3 band does not cross any longer, an effect that is further enhanced by . Second, the La-5 band at M, which also undergoes a large negative correction of almost 1Β eV, now stays above the Fermi level by eV. In fact, it only reaches the upper Ni-3 band without crossing it. We can see that, with respect to the band positions at M, the epitaxial strained configuration is intermediate between 29.5Β GPa and ambient pressure. Although our biaxial strain simulation implements a non-hydrostatic realization, we can compare it to the hydrostatic pressure case where the La-5 BOC is degenerate with Ni-3 BOC at M. Following this analogy, from Fig.Β 4b, where we have also reported the band positions for the thin film, we can tentatively attribute our epitaxial strain simulation as equivalent to the 14Β GPa hydrostatic pressure case, that is precisely at the superconducting onset. This could be an indication that the relative position and overlap of BOC at M, might play a role as a triggering mechanism of superconductivity.
ARPES spectra have been measured on 1, 2 and 3 unit-cell epitaxial La2.85Pr0.15Ni2O7 films grown on SrLaAlO4 substrates [57]. It is claimed that, in addition to the and sheets, the pocket is also present at the Fermi level. In another more recent ARPES experiment on a La2PrNi2O7 thin film [73], the pocket is absent from the Fermi level and the Ni-3 band is found 70Β meV below the Fermi energy, which compares more favorably with our result of 200Β meV. In any case, the comparison with our electronic structure is difficult because they considered a Pr doped system. From our side, we did not include the neighborhood effect of the substrate which is certainly affecting the electronic structure of such thin films. Our results should then be considered as an ab initio prediction of the electronic structure which could be measured for a La3Ni2O7 thin film of enough large thickness. Or, possibly, for a non-hydrostatic planar strain applied on a bulk sample that can be realized experimentally by applying a tensile strain to the axis, which will induce a biaxial compressive strain of the plane by Poisson law.
We can conclude that our calculation on the epitaxial strained case presents an electronic structure very similar to the bulk at 14Β GPa pressure. Correlation effects are very important on both the lower Ni-3 band, causing the disappearance of the pocket, and on the increased relevance of the La-5 band at low energy.
IV Conclusions
In conclusion, we have studied pressure and biaxial strain effects on the electronic structure of La3Ni2O7, taking into account correlation effects within the ab initio approximation to the self-energy. We confirm that the electronic structure of La3Ni2O7 is extremely sensitive to the internal atomic positions of its crystal structure. correlation effects shift downward the lower Ni-3 band, so to remove the hole pocket from the Fermi level without introducing any adjustable parameter. Therefore, the Fermi surface is composed only of the cuprate-shape sheet plus the nickelate-specific cylinder , both showing an effective 1D character, and is in good agreement with ARPES experiments [56] which did not find any pocket. Additionally, we observe a non-negligible drop in the quasiparticle spectral weight on the Ni-3 states, both and , so to configure a little departure from a Fermi liquid behavior. Finally, we have shown that not only correlations, but also pressure or biaxial strain play a crucial role in pulling down the La-5 band at and M towards the Fermi level, which may induce self-doping. In any case, at pressures and strains of interest of superconductivity, this band might play an important role at low energy, a fact which is often overlooked in effective-model studies.
V Acknowledgments
We thank Markus Holzmann for useful discussions. We acknowledge financing from the LABEX LANEF project NICOS. Computing time has been provided by French GENCI and Grenoble CIGRID.
Appendix A Crystal structures used in the calculations
TAB.Β 1 reports the crystal structures in the conventional cell used for the calculation at 29.5Β GPa and for the in-plane strained system.
In these two cases the parameters were chosen with a trade off between the values from the literature and obtained via a structural optimization.
TAB.Β 2 reports the crystal structures used for the extra pressure calculations, all obtained via structural optimizations.
| 29.5 GPa [1, 24] | a=3.7148 | c=19.7340 | ||
| Ni | 0 | 0 | 0.096 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.095 | |
| O2 | 0 | 0 | 0.204 | |
| O3 | 0 | 0 | 0 | |
| epitaxial strained (thin film at ambient pressure) [5] | a=3.77 | c=20.6927 | ||
| Ni | 0 | 0 | 0.095 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.096 | |
| O2 | 0 | 0 | 0.205 | |
| O3 | 0 | 0 | 0 | |
| 0 GPa | a=3.8408 | c=20.2929 | ||
| Ni | 0 | 0 | 0.097 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.096 | |
| O2 | 0 | 0 | 0.204 | |
| O3 | 0 | 0 | 0 | |
| 14 GPa | a=3.7533 | c=19.7621 | ||
| Ni | 0 | 0 | 0.098 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.096 | |
| O2 | 0 | 0 | 0.204 | |
| O3 | 0 | 0 | 0 | |
| 40 GPa | a=3.6378 | c=19.1413 | ||
| Ni | 0 | 0 | 0.098 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.095 | |
| O2 | 0 | 0 | 0.204 | |
| O3 | 0 | 0 | 0 | |
Appendix B Sensitiveness of the electronic structure to internal atomic positions
In this Section we will show the critical dependence of the electronic structure from the crystal structure. In particular, tiny displacements of the internal atomic coordinates can strongly affect the resulting electronic structure. The latter is then further modified by the variation of the lattice parameters under pressure which is fundamental to trigger superconductivity in La3Ni2O7.
In our first attempt to calculate the DFT and the electronic structures of the bilayer at 29.5 GPa, we used the experimental atomic positions determined by X-ray diffraction (XRD) given in the extended data table I in Sun et al.Β [1] for the Fmmm phase and adapted to the as also done in Sakakibara et al.Β [24]
The resulting DFT and Fermi surfaces (Fig.Β 6) show the appearance of a cylindrical Fermi sheet centered on which is not observed in the ARPES experiment of Yang et al.Β [56] This fake Fermi sheet originates from the crossing at the Fermi level of the La-5 character band, as one can see in Fig.Β 7a for PBE and in Fig.Β 7b for . It has to be considered as an artifact due to a non correct configuration of the internal atomic positions. In XRD the accuracy on the internal atomic positions is in general lower than on the lattice parameters, especially concerning the oxygen atoms. Due to this drawback we decided for this work to rely on theoretical DFT-PBE relaxed internal atomic positions (see Tab. 1), as explained in Sec. II.2 of the main text.
The DFT bandplot of this relaxed structure is shown on Fig.Β 7c and has to be compared with Fig.Β 7a corresponding to the experimental structure. In particular, we remark that the self-doping induced by the La-5 band is absent for the relaxed structure. The Ni- bands are less affected by these internal atomic positions differences, but quantitative discrepancies are visible, for example the size of the Fermi sheet. Some of these effects were already previously reported [45].
Since any real or fake modification of the electronic structure at the Fermi level may have important consequences for superconductivity, we tried to identify what precisely triggers the La-5 band lowering. To this purpose we computed (at the DFT level only) the bandplot of a structure similar to the relaxed one, but with the La atoms in the spacer (La1 in Tab. 1) artificially shifted away from the NiO2 plane by 0.14Β Γ , which corresponds to 0.7% of (see Fig. 8). The associated bandplot is shown on Fig. 7d and reproduces well the self-doping behavior of the rare-earth band that is observed on the DFT bandplot of Fig.Β 7a, and to a lesser extent on the bandplot of Fig.Β 1d. We could have rather moved the apical oxygens of the spacer toward the NiO2 planes, but this would have changed the shape of the octahedron and potentially the crystal field splitting. In order to keep the ordering of the Ni bands, we have preferred to move the lanthanum of the spacer towards the apical oxygens, which should be equivalent.
What is described here is a large effect that depends on tiny shifts along the direction of the internal atomic positions which do not imply any symmetry breaking. Usually, atomic shifts of such tiny amplitude and without symmetry breakdown, should not affect the electronic structure with such large qualitative modifications. This is a first important indication that the electronic structure of La3Ni2O7 results from a delicate balance of many factors, already starting from internal atomic positions. Tiny atomic displacements, further enhanced by correlations, can provide self-doping and new Fermi surface pockets, bringing to a completely different Fermi level scenario.
We hope that future experiments, for example using other techniques like neutron diffraction measures, could be helpful in clarifying the internal atomic positions which are critical for the electronic structure.
We anticipate that and we are working on a more complete study [74] about this effect and its correct description, both the internal atomic positions and the numerical methods employed being important. This can further help shading light on the aspects which could be relevant to superconductivity in RP nickelates.
Appendix C Fermi energy calculation: standard bisection vs DOS integration methods
The indirect gap of 30 meV between the two incipient bands found in the calculation done at 29.5 GPa is beyond the accuracy by which the Fermi level can be calculated. This is already a consequence of the finite smearing temperature used to calculated metals, of the order of Β Ha Β meV and up to Β Ha Β meV (the Abinit default is Β Ha, while it is half that in Quantum Espresso). We must also keep in mind that any change in the parameters used to generate the pseudo-potentials can affect DFT Kohn-Sham energies by even 100200Β meV. This is also the reason why the quasiparticle energy convergence is never pushed beyond 100Β meV. A further indetermination of the Fermi energy arise in particular from the choice of the method used to calculate it. In this Section we present the comparison between two of them: the classical method using a bisection algorithm, as implemented in Abinit or Quantum Espresso; and the Wannier interpolation as implemented in Wannier90, followed by the integration of the total DOS done with our own Python code using the Simpson method.
In this Section we show how critical the calculation of the Fermi energy can be on the electronic structure and the Fermi surface of this system. For this reason, we take the most critical case of 29.5Β GPa and we check the two methods here above described (for short labeled the bisection and DOS methods), keeping as close as possible to their defaults when there is matter of choice. In Fig.Β 9 we show a zoom around the Fermi level of the bandplot, as in Fig.Β 1d but reporting the different Fermi energies calculated using a smearing temperature of Ha, with the corresponding Fermi surfaces below. First of all, one can remark the robustness of the and sheets with respect to a large variation of (120 meV), which is due to the more dispersive character of these bands compared with to two incipient ones.
With the bisection method ( traced by a black line and set as the origin), a pocket of electrons () appears around the M point in the plane (see BZ scheme on Fig.Β 1c), due to the La-5 band crossing. On the other hand, with the DOS method and using the least arbitrary Wannier functions set among all possible sets, the PWF one, the Fermi energy is 87 meV (purple line) lower. In this case, there is no pocket of electrons at M and, in contrast, a pocket of holes is opened in the XP direction in the calculation, and the Fermi surface becomes again very similar to the DFT-PBE one, i.e.Β with the reappearance of the pocket in the BZ corners. As it can be seen, different methods provide qualitatively different Fermi surfaces leading to different low energy physics.
In order to improve our estimate on the Fermi energy, we have performed a full DFT+ calculation reducing the smearing temperature to Ha Β meV. The corresponding quasiparticle energies are indicated by pink stars and the associated by the green dashed line in Fig.Β 9. As we can see, now the Fermi level situated more or less in the middle of the indirect band gap, at meV compare to the calculated with Ha. As a consequence both the electron pocket and the hole pocket are suppressed. And this is the Fermi level we have used to plot the Fermi surface in Fig.Β 1b.
In practice, in a system with such a delicate balance, any defect, vacancy (La3Ni2O7 is usually oxygen deficient) or impurity, introduces a tiny perturbation in the Fermi energy which will open the electron or the hole pocket. In any case, the Fermi surface showed on Fig.Β 1b corresponds to a theoretical ideal situation. Therefore, we expect that a real experiment will probably not measure the ideal situation. Therefore we leave a margin of error of 50Β meV on the Fermi level to keep into account both experimental and theoretical uncertainties.
Notice that 29.5Β GPa is a really critical pressure for the La3Ni2O7 electronic structure. Indeed, at ambient pressure the La-5 band is almost 0.5Β eV above its position at 29.5Β GPa (see Fig.Β 4b), and the bisection method set the Fermi energy well below the bottom of this band, with no possibility of an electron pocket opened at M. Consequently, the Fermi surface would be exactly that in Fig.Β 1b, with the stable and sheets and no or pockets at all, from ambient pressure on, and without any relevant modification up to 29.5Β GPa.
Appendix D Orbital character projection: pure atomic orbitals vs Wannier functions
Fig. 10 shows three DFT PBE band character plots of La3Ni2O7 at 29.5 GPa: the first one using projections on (pseudo) atomic orbitals calculated with Quantum Espresso, which we will consider as the reference; and the two others with Wannier90 but using different WF sets, namely projected Wannier functions (PWF) and maximally-localized Wannier functions (MLWF). PWF are the Wannier90 first guess and correspond to WF with imposed well defined orbital character (in the input file), though not purely hydrogenoid atomic orbitals [75, 76] (see below). Therefore, PWF is the set of WF one obtains from Wannier90 by setting to zero the number of iterations in the Wannierisation procedure (not in the disentanglement one). PWF might not be the set best suited to a smooth interpolation of bands. However, in this case we used 67 WF, which is a set large enough to allows an already smooth interpolation. On the other hand, MLWF are obtained after a number of Wannierisation iterations sufficient to minimize the Marzari-Vanderbilt functional towards the global or a local minimum, which correspond to a maximum in the localization of the WF. MLWF can loose their original orbital character, and they can even localize outside atoms and in the interstitial, if the target is to have the most localized WF.
The Quantum Espresso and the PWF band plot characters are very consistent between themselves, whereas the MLWF shows important discrepancies for the La- orbital character which lose weight on the band of interest and is more hybridized with the 4 manifold. It must be kept into account that the projection onto MLWF, beyond to strongly depend on the chosen set of MLWF (e.g.Β localized on atoms, on bonds, or on anti-bonding positions), is different from the projection onto atomic wave functions. But there is large arbitrariness also for atomic wave functions. For example whether projecting onto all-electron wave functions (as in all-electron codes), or pseudo-atomic wave functions (e.g.Β Quantum Espresso and Abinit, with some minor differences between them), or spherical harmonics (e.g.Β VASP). In conclusion, projections can be a convenient label to identify bands, establishing a convention. However, any physics argument built on them is always risky.
The comparison done here at the DFT level validates the use of PWF to show the orbital characters on the band character plots at the level (Fig. 1d and 5), for which a dense orbital character projection on bands is not possible because the energies are only sampled on a grid. Note that this orbital character comparison between Quantum Espresso and Wannier90 has been done for the PDOS as well, providing the same conclusion.
Appendix E Additional orbital character projections
On Fig.Β 12 we provide more precise band character plots than on Fig.Β 2b, giving details on the specific orbitals involved in the Bloch states. We distinguish the two spacer La atoms from the inner one lying in between the two NiO2 planes; as well as the planar O atoms lying in the NiO2 planes, and the apical O atoms of the spacer (ap.Β sp.) or of the inner layer (ap.Β inner). See Fig.Β 13 which shows the atoms labeling on the Vesta structure. For O-2, we also distinguish and orbitals.
We remark that there is a non-negligible La-4 orbital character on the lowered La-5 band which almost reaches the Fermi level. We also remark that in-plane O-2 orbitals have a strong contribution on what we label Ni-3 bands, especially above the Fermi level.
Appendix F Convergence in the number of bands
In a calculation, the number of bands taken into account in the calculation of the self-energy is critical. A convergence study is therefore needed. Moreover, this study is motivated by the discrepancy on the Ni-3 upper and lower bands which we have found between our results and the results of Christiansson et al.Β [45] from one side, and of You et al.Β [15] from another, especially regarding the shift the latter obtained on the upper Ni-3 band from DFT to . In our original calculation, we used 225 bands. For this test we rise up to 300 for the calculation of energies of three bands around the Fermi level. The energies obtained are plotted on top of the original calculation in Fig.Β 14. As we can see, the original calculation is already enough converged, the maximal difference for the set of checked bands being of Β meV, validating the electronic structure presented in the main text.
Appendix G Fermi surfaces of the substrate constrained and ambient pressure cases
Appendix H Bloch state plots
In this section we show some selected Bloch states that are interesting to visualize in real space. Notice that these Bloch states are calculated in DFT PBE, as our one-shot first order perturbation theory calculation does not recalculate the wavefunctions, but only corrects the energies.
The lower and upper Ni-3 bands are sometimes respectively labeled as bonding and anti-bonding, according to a chemical convention. For these two states we plot in Fig.Β 17 the iso-surfaces of the Bloch wave functions at the X -point which lies in the corner of the BZ. As we can see, these names are not appropriated, the upper state being more βbondingβ than the lower one, with more weight on orbitals of inner oxygen atoms lying in between NiO2 planes.
Another Bloch state worth to be shown is the one corresponding to the La-5 band, which is the most sensitive to pressure, and in particular at the M point where it reaches its minimum. In Fig.Β 18 we see that these electrons are delocalized in the spacer, which explain the highly dispersive shape of the corresponding band. From 0 to 29.5Β GPa, this state is not occupied. It starts to be occupied beyond 29.5Β GPa, inducing self-doping of the NiO2 planes.

### v2

@@
 Abstract
-The recent discovery of superconductivity in La3Ni2O7 at a critical temperature above 80Β K points to a non-conventional pairing mechanism in nickelates as in cuprates, possibly due to electronic correlations.
+The recent discovery of superconductivity in La3Ni2O7 at a critical temperature above 80 K points to a nonconventional pairing mechanism in nickelates as in cuprates, possibly due to electronic correlations.
 We have calculated from first principles the electronic structure of La3Ni2O7 under the effect of pressure and epitaxial strain including correlations by the approximation to the many-body self-energy.
-We find that the Fermi surface is composed of a characteristic cuprate-shape sheet plus a nickelate-specific cylinder , both from Ni orbitals, with a non-negligible drop in the quasiparticle weight and an effective 1D character.
+We find that the Fermi surface is composed of a characteristic cuprate-shape sheet plus a nickelate-specific cylinder , both from Ni orbitals, with a nonnegligible drop in the quasiparticle weight and an effective 1D character.
 This topology results from a delicate balance between the Ni-3 hole pocket , which is suppressed by correlations, and an emerging La-5 electron pocket induced by both correlation and pressure/strain effects and whose role at low energy has been neglected so far.
@@
 The recent discovery of superconductivity (SC) in Ruddlesden-Popper (RP) nickelates [1, 2, 3] has introduced a new direction in the field of high-temperature superconductivity.
-Notably, superconductivity has been observed in bulk samples of the bilayer La3Ni2O7 under high pressure [1], with the superconducting phase forming a right-triangular region spanning 14 to 80 GPa, and reaching a maximum onset critical temperature of approximately 80Β K at 18 GPa [4].
-More recently, superconductivity has also been reported at Β K in La3Ni2O7 thin films under epitaxial compressive strain at ambient pressure [5, 6].
-While the superconducting mechanism is still debated in infinite-layer LnNiO2 nickelates [7, 8, 9, 10, 11], where critical temperatures do not overstep 30Β K [12, 13], the high critical temperature observed in La3Ni2O7 is difficult to reconcile with a BCS electron-phonon pairing mechanism [14, 15], thus pointing to unconventional superconductivity like in cuprates.
-Indeed, cuprates and nickelates share many analogies: they are both organized in a layered crystal structure with CuO2 or NiO2 planes; and present the same βcuprate-shapeβ Fermi surface sheet of main 3 character, although nickelates bilayer and trilayer show a strong Ni-3 hybridization in the antinodal direction.
-However, there are also important differences: the constant presence of an anti-ferromagnetic insulating parent phase in cuprates has no analogous in nickelates.
+Notably, superconductivity has been observed in bulk samples of the bilayer La3Ni2O7 under high pressure [1], with the superconducting phase forming a right-triangular region spanning 14 to 80 GPa, and reaching a maximum onset critical temperature of approximately 80 K at 18 GPa [4].
+More recently, superconductivity has also been reported at K in La3Ni2O7 thin films under epitaxial compressive strain at ambient pressure [5, 6].
+While the superconducting mechanism is still debated in infinite-layer LnNiO2 nickelates [7, 8, 9, 10, 11], where critical temperatures do not overstep 30 K [12, 13], the high critical temperature observed in La3Ni2O7 is difficult to reconcile with a BCS electron-phonon pairing mechanism [14, 15], thus pointing to unconventional superconductivity like in cuprates.
+Indeed, cuprates and nickelates share many analogies: they are both organized in a layered crystal structure with CuO2 or NiO2 planes; and present the same “cuprate-shape” Fermi surface sheet of main 3 character, although nickelates bilayer and trilayer show a strong Ni-3 hybridization in the antinodal direction.
+However, there are also important differences: the constant presence of an antiferromagnetic insulating parent phase in cuprates has no analogous in nickelates.
 For what concerns the electronic structure, the fact that in nickelates the Ni-3 states are closer to the Fermi level than in cuprates, gives rise to at least another Fermi sheet and possibly to a multiband character superconductivity.
@@
 In this work we study from ab initio the electronic structure of La3Ni2O7 including correlations at the level of the approximation to the self-energy [59, 60, 61, 62].
-In contrast to the Kohn-Sham eigenvalues used in density-functional theory (DFT), which are the energies of a fictitious non-interacting system, quasiparticle energies have a direct physical interpretation as the excitation energies measured in ARPES.
-The approximation takes into account electronic exchange and correlations without introducing any adjustable parameter, neither in the DFT functional (e.g.Β an hybrid mixing parameter) nor in the Hamiltonian (e.g.Β a Hubbard interaction term).
+In contrast to the Kohn-Sham eigenvalues used in density-functional theory (DFT), which are the energies of a fictitious noninteracting system, quasiparticle energies have a direct physical interpretation as the excitation energies measured in ARPES.
+The approximation takes into account electronic exchange and correlations without introducing any adjustable parameter, neither in the DFT functional (e.g.
+an hybrid mixing parameter) nor in the Hamiltonian (e.g.
+a Hubbard interaction term).
 The weight of correlation effects and their importance in this system is revealed by a comparison between the electronic structure and the DFT band plot.
+However, in this work we applied the one-iteration approach which can keep some reminiscence of the DFT starting point.
 We also investigate the electronic structure evolution of bulk La3Ni2O7 with respect to pressure, as well as in presence of in-plane constraints that simulate a thin film grown on a substrate.
 Our purpose is to highlight aspects which might be relevant to explain superconductivity that in this system is induced by pressure without doping.
-The focus is at the Fermi level but also at higher energy features which are less affected by the estimated 0.1Β eV best absolute accuracy, and which might be relevant in the competition between different pairing symmetries according to some proposed superconducting mechanisms [24].
+The focus is at the Fermi level but also at higher energy features which are less affected by the estimated 0.1 eV best absolute accuracy, and which might be relevant in the competition between different pairing symmetries according to some proposed superconducting mechanisms [24].
 Our work describes some of the electronic structure features expected to be relevant by theoretical models so far proposed [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and provides also indications for future works.
-The paper is organized as follows: after introducing our methods in Sec.Β II, we will first compare our calculations with previous ab initio works done at 29.5Β GPa with the crystal structure used by almost all DFT as well as two calculations [45, 15] (Sec.Β III.1).
-We will then show the evolution with pressure of the electronic structure for the bulk crystal in Sec.Β III.2, and finally introduce in Sec.Β III.3 our results for the structure with epitaxial strain of 1.8%.
+The paper is organized as follows: after introducing our methods in Sec.
+II, we will first compare our calculations with previous ab initio works done at 29.5 GPa with the crystal structure used by almost all DFT as well as two calculations [45, 15] (Sec.
+III.1).
+We will then show the evolution with pressure of the electronic structure for the bulk crystal in Sec.
+III.2, and finally introduce in Sec.
+III.3 our results for the structure with epitaxial strain of 1.8%.
 Our results will be compared with the available experimental data from photoemission (ARPES) experiments.
@@
 Our methodology relies as a first step on density-functional theory (DFT) calculations in the PBE approximation [63], using the plane waves code Abinit [64].
-We use norm-conserving pseudo-potentials from the PseudoDojo table library [65].
-The La pseudo-potential takes 4 electrons in conduction and does not freeze them in the core.
-We have used a 161616 standard sampling of the Brillouin zone (BZ) and a plane waves (PW) cutoff of 65Β Ha for the DFT-PBE starting calculation, as well as a Gaussian smearing of 0.01 Hartree.
+We use norm-conserving pseudopotentials from the PseudoDojo table library [65].
+The La pseudopotential takes 4 electrons in conduction and does not freeze them in the core.
+We have used a 161616 standard sampling of the Brillouin zone (BZ) and a plane waves (PW) cutoff of 65 Ha for the DFT-PBE starting calculation, as well as a Gaussian smearing of 0.01 Hartree.
 In the subsequent calculations, the BZ sampling has been reduced to 666, whereas the PW cutoffs have been reduced to 40 Ha for the representation of the wave functions and for the exchange self-energy, and further down to 15 Ha for the correlation self-energy and the screening.
-We included 187 bands for the calculation of the screening and 225 for the self-energy (see Sect.Β F in the AppendixΒ for convergence studies).
-We have carried just only one iteration (i.e.Β ) on top of PBE, using a Godby-Needs plasmon-pole model [66] and a shift of 0.1Β eV to avoid poles/divergences.
+We included 187 bands for the calculation of the screening and 225 for the self-energy (see Sect.
+G in the Appendix for convergence studies).
+We have carried just only one iteration (i.e.
+) on top of PBE, using a Godby-Needs plasmon-pole model [66] and a shift of 0.1 eV to avoid poles/divergences.
 The final band plots, total and projected density of states (DOS and PDOS), as well as Fermi surfaces, were interpolated from the 666 to a denser -mesh by the Wannier90 code using a set of 67 projected Wannier functions (PWFs) with the closest possible character to atomic orbitals, namely O-, Ni-, La- and La-.
-The band orbital characters and PDOS identified by these PWFs are almost coincident with what one would obtain by projecting Bloch wave functions directly onto real atomic orbitals, which is not the case when selecting a set of maximally-localized Wannier functions (MLWFs) (see App.Β D).
-In the bandplots showing projected orbital character, the area of the dots scales as the orbital contribution squared, so that the width of the line is proportional to it.
-All calculations have been done with a Gaussian smearing temperature of 0.01Β Ha to ease self-consistency convergence and optical absorption onset stability (which in any case has little relevance on main plasmons, energy-loss and screening ).
-For the most critical case at Β GPa we have checked that a calculation with enforced at Β Ha in both DFT and calculations provides the same physical picture, and in particular the same Fermi surface as in Fig.Β 1b, with only minor adjustments of less than Β meV on band positions, mostly due to the fact that the chemical potential at larger temperature is a less accurate estimate for the Fermi energy (see App.Β C).
+The band orbital characters and PDOS identified by these PWFs are almost coincident with what one would obtain by projecting Bloch wave functions directly onto real (pseudo) atomic orbitals, as we have checked using the Quantum Espresso code with equivalent parameters, which is not the case when selecting a set of maximally localized Wannier functions (MLWFs) (see App.
+E).
+In the bandplots showing projected orbital character, the width of the dots is proportional to the orbital contribution.
+All calculations have been done with a Gaussian smearing temperature of 0.01 Ha to ease self-consistency convergence and optical absorption onset stability (which in any case has little relevance on main plasmons, energy-loss and screening ).
+For the most critical case at GPa we have checked that a calculation with enforced at Ha in both DFT and calculations provides the same physical picture, and in particular the same Fermi surface as in Fig.
+1b, with only minor adjustments of less than meV on band positions, mostly due to the fact that the chemical potential at larger temperature is a less accurate estimate for the Fermi energy (see App.
+D).
 II.2 Crystal structure and Brillouin zone
-To determine the crystal structure of La3Ni2O7 at ambient pressure, and then in the superconducting range at 14, 29.5 and 40 GPa, we have performed DFT PBE structural relaxations which provided the I4/mmm space group as the lowest energy structure at all considered pressures, except at 0Β GPa where the Amam structure is more stable.
-Nevertheless, our calculations have all referred to the I4/mmm structure even at ambient pressure to ease the comparison and better follow the evolution of the electronic structure with pressure, from the superconducting range down to 0Β GPa.
-For all the cases we have used the relaxed lattice parameters and the internal atomic position out of our PBE relaxations, except for the case at 29.5Β GPa which we treated separately.
-Indeed, 29.5Β GPa is the most studied case in the theoretical literature, so that we use it as a benchmark case to compare with previous calculations.
+To determine the crystal structure of La3Ni2O7 at ambient pressure, and then in the superconducting range at 14, 29.5 and 40 GPa, we have performed DFT PBE structural relaxations which provided the I4/mmm space group as the lowest energy structure at all considered pressures, except at 0 GPa where the Amam structure is more stable.
+Nevertheless, our calculations have all referred to the I4/mmm structure even at ambient pressure to ease the comparison and better follow the evolution of the electronic structure with pressure, from the superconducting range down to 0 GPa.
+For all the cases we have used the relaxed lattice parameters and the internal atomic position out of our PBE relaxations, except for the case at 29.5 GPa which we treated separately.
+Indeed, 29.5 GPa is the most studied case in the theoretical literature, so that we use it as a benchmark case to compare with previous calculations.
 Thus, for the sake of comparison, we decided to stick to the lattice parameters and atomic positions used in most of previous works [45, 24].
-Therefore, we consider a tetragonal structure with Β Γ Β and Β Γ Β as Sakakibara et al.Β [24].
-In fact, instead of the initial space-group symmetry proposed for La3Ni2O7 at 29.5Β GPa, our calculations predict the space-group symmetry with lattice parameters that are in good agreement with the experimental ones [67].
-Further, in Ref.Β 45, it was noticed that the electronic structure of La3Ni2O7 at 29.5Β GPa is extremely sensitive to internal atomic positions and that the experimental ones reported by Sun et al.Β from the X-ray diffraction [1] likely yield artifacts at the Fermi level.
-We provide a detailed analysis of this question in AppendixΒ B.
-To simulate epitaxial strain, we enforced the tetragonal lattice parameter to the experimentally observed value of 3.77Β Γ Β [5] and let the system relax in the direction, which provided Β Γ .
-This crystal structure corresponds to a bulk non-hydrostatic planar stressed system, or equivalently a system pulled along the axis, that we will use to simulate an epitaxially strained thin film.
-All our crystal structures are detailed in App.Β A.
-All our electronic structures are reported in the body-centered tetragonal (BCT) Brillouin zone (BZ) which we report in Fig.Β 1c by black lines, while the blue line shows the path followed in our band plots with -point labels using the convention of Hinuma et al.
-Β [68].
+Therefore, we consider a tetragonal structure with Å and Å as Sakakibara et al.
+[24].
+In fact, instead of the initial space-group symmetry proposed for La3Ni2O7 at 29.5 GPa, our calculations predict the space-group symmetry with lattice parameters that are in good agreement with the experimental ones [67].
+Further, in Ref.
+45, it was noticed that the electronic structure of La3Ni2O7 at 29.5 GPa is extremely sensitive to internal atomic positions and that the experimental ones reported by Sun et al.
+from the x-ray diffraction [1] likely yield artifacts at the Fermi level.
+We provide a detailed analysis of this question in App.
+B.
+To simulate epitaxial strain, we enforced the tetragonal lattice parameter to the experimentally observed value of 3.77 Å [5] and let the system relax in the direction, which provided Å.
+This crystal structure corresponds to a bulk nonhydrostatic planar stressed system, or equivalently a system pulled along the axis, that we will use to simulate an epitaxially strained thin film.
+All our crystal structures are detailed in App.
+A.
+All our electronic structures are reported in the body-centered tetragonal (BCT) Brillouin zone (BZ) which we report in Fig.
+1c by black lines, while the blue line shows the path followed in our band plots with -point labels using the convention of Hinuma et al.
+[68].
 Notice that in this convention, the X and M -points do not correspond to the canonical ones used in the 2D BZ of the square lattice widely used in the model literature: X is at the BZ corner in the nodal direction and M lies in the plane.
 We also indicate in grey the BZ corresponding to the conventional cell, namely simple tetragonal, which ease the comparison with 2D BZs.
-In the same Fig.Β 1c, we report also the simple orthorhombic BZ and associated -path to compare with the ARPES experiment [56] which referred to the tilted Amam structure needing a supercell compared to ours; and finally the face-centered orthorhombic BZ (and associated -path) to compare with the calculation of Christiansson et al.Β [45] who referred to the crystal structure.
+In the same Fig.
+1c, we report also the simple orthorhombic BZ and associated -path to compare with the ARPES experiment [56] which referred to the tilted Amam structure needing a supercell compared to ours; and finally the face-centered orthorhombic BZ (and associated -path) to compare with the calculation of Christiansson et al.
+[45] who referred to the crystal structure.
 III Results
 III.1 Electronic structure at 29.5 GPa
-We start discussing the electronic structure of La3Ni2O7 at the pressure of 29.5Β GPa over which most of previous theoretical calculations have been performed.
-Our PBE and Fermi surfaces are shown in Fig.Β 1a and 1b, while Fig.Β 1d shows the corresponding band structures.
+We start discussing the electronic structure of La3Ni2O7 at the pressure of 29.5 GPa over which most of previous theoretical calculations have been performed.
+Our PBE and Fermi surfaces are shown in Fig.
+1a and 1b, while Fig.
+1d shows the corresponding band structures.
 In the same figure, selected orbital character contributions are plotted on the band structure, as well as the projected densities of states (PDOS) and the total density of states (DOS).
-We briefly mention that, like it was already found [69, 7, 8] in infinite-layer nickelates, e.g.Β LaNiO2, correlations shift 2Β eV up the flat bands with predominant La-4 character, moving them further away from the Fermi level and out of the low-energy region, and shift 1Β eV down the O-2 manifold, leading to an increased ionic charge-transfer energy between Ni-3 and O-2 states.
+We briefly mention that, like it was already found [69, 7, 8] in infinite-layer nickelates, e.g.
+LaNiO2, correlations shift 2 eV up the flat bands with predominant La-4 character, moving them further away from the Fermi level and out of the low-energy region, and shift 1 eV down the O-2 manifold, leading to an increased ionic charge-transfer energy between Ni-3 and O-2 states.
 These aspects were previously detailed [69] and will not be again discussed here where we focus on the low-energy region.
@@
 The sheet is a large cylindrical sheet centered around ; is the sheet with the characteristic shape of the cuprate Fermi surface; and is a small hole pocket located at the Brillouin zone (BZ) corners which is highly dispersive along .
-The and sheets are primarily of character, though they show significant hybridization with orbitals in the anti-nodal direction, as well as with O-2 states (see Fig.Β 2b).
+The and sheets are primarily of character, though they show significant hybridization with orbitals in the antinodal direction, as well as with O-2 states (see Fig.
+2b).
 In contrast, the pocket has an almost pure character and is consistently present in PBE (or LDA) calculations.
-The Fermi surface, in contrast, has no hole-pockets, as shown in Fig.Β 1b.
-This result is totally consistent with experimental ARPES data [56], with respect to which we provide a detailed comparison in Sec.Β III.2 below, since it was performed at ambient pressure.
-This difference between DFT and is due to a slight downward shift (7080 meV) of the lower Ni-3 band along the XP direction (see Fig.Β 1d).
+The Fermi surface, in contrast, has no hole-pockets, as shown in Fig.
+1b.
+This result is totally consistent with experimental ARPES data [56], with respect to which we provide a detailed comparison in Sec.
+III.2 below, since it was performed at ambient pressure.
+This difference between DFT and is due to a slight downward shift (7080 meV) of the lower Ni-3 band along the XP direction (see Fig.
+1d).
 The absence of the pocket in experiment has been discussed also in the literature, and several semi-empirical corrections have been shown to reconcile theoretical calculations with the ARPES experiment.
-These methods either include the use of a weighted Fock exchange operator, as by HSE [54] or any other hybrid approach; or place more strength on correlation by introducing a Hubbard repulsion term of Β eV on Ni-3 electrons, like it was done in the calculation of Yang et al.Β [56], with the effect to lower this band.
-On the basis of this fact, Yang et al.Β [56] concluded that La3Ni2O7 is a strongly correlated system.
+These methods either include the use of a weighted Fock exchange operator, as by HSE [54] or any other hybrid approach; or place more strength on correlation by introducing a Hubbard repulsion term of eV on Ni-3 electrons, like it was done in the calculation of Yang et al.
+[56], with the effect to lower this band.
+On the basis of this fact, Yang et al.
+[56] concluded that La3Ni2O7 is a strongly correlated system.
 Here we show that, a calculation using physical quasiparticle energies out of a many-body self-energy, even in a non self-consistent approximation, is enough to get rid of the sheet at the Fermi level and achieve good agreement with the experiment.
-Our results are in line with the findings of Ryee et al.Β [55] who excluded both a transition to a Mott insulator or bad metal scenario.
-The disappearance of the hole pocket has also been reported in a previous work [45] (see Fig.Β 1c for -point correspondence).
+Our results are in line with the findings of Ryee et al.
+[55] who excluded both a transition to a Mott insulator or bad metal scenario.
+The disappearance of the hole pocket has also been reported in a previous work [45] (see Fig.
+1c for -point correspondence).
 They performed a calculation based on a reduced four-orbital tight-binding model restricted to the Ni- manifold.
-This model does not include the La-5 band, which, as we will discuss in the next section, plays a non-negligible role in the low energy spectrum.
+This model does not include the La-5 band, which, as we will discuss in the next section, plays a nonnegligible role in the low energy spectrum.
 Their model also employs an effective screening that is renormalized to account for the missing degrees of freedom, and refers to the Fmmm structural phase.
 Despite these significant methodological differences, our results are in remarkable agreement with theirs regarding the absence of the pocket.
-However, Christiansson et al.Β do not further comment on this disappearance in their calculation, because their subsequent EDMFT simulationsβperformed on top of both DFT and βreintroduce and even enhance the spectral weight of the lower Ni-3 band at the Fermi level.
-On this basis, they conclude that βthe dominance of the orbital distinguishes La3Ni2O7 from the infinite-layer nickelates or the cuprates,β which are instead dominated by character.
-This interpretation is not supported by our results (see Sec.Β III.1.4 for the spectral weights), which show a predominantly character at low energy.
-Our findings rather align with the recent calculations of You et al.Β [15].
-Interestingly, Ryee et al.Β [55] found the pocket in both DFT and dynamical mean-field theory (DMFT) calculations, but not in cluster-DMFT (CDMFT) which is therefore more in agreement with the finding.
-This is an indication that non-local correlations, as provided by a CDMFT or a -dependent self-energy, are important to describe such modification of the Fermi surface topology.
+However, Christiansson et al.
+do not further comment on this disappearance in their calculation, because their subsequent EDMFT simulations—performed on top of both DFT and —reintroduce and even enhance the spectral weight of the lower Ni-3 band at the Fermi level.
+On this basis, they conclude that “the dominance of the orbital distinguishes La3Ni2O7 from the infinite-layer nickelates or the cuprates,” which are instead dominated by character.
+This interpretation is not supported by our results (see Sec.
+III.1.4 for the spectral weights), which show a predominantly character at low energy.
+Our findings rather align with the recent calculations of You et al.
+[15].
+Interestingly, Ryee et al.
+[55] found the pocket in both DFT and dynamical mean-field theory (DMFT) calculations, but not in cluster-DMFT (CDMFT) which is therefore more in agreement with the finding.
+This is an indication that nonlocal correlations, as provided by a CDMFT or a -dependent self-energy, are important to describe such modification of the Fermi surface topology.
 However, CDMFT also find an important modification to the cuprate-shape sheet which loses most of its 1D character (see below).
@@
 III.1.2 La-5 band lowering and self-doping
-Beside this important qualitative modification of the Fermi surface, correlation corrections induce an even more evident and quantitative change in the bandwidth of the 5 band contributed by the two La spacer atoms only (i.e., with no contribution by the La inner atom lying in between the NiO planes, see Fig.Β 13).
-This La-5 band, which in DFT-PBE is well above the Ni-3 one, is lowered by corrections, as it can be seen at in Fig.Β 1d.
-At the M point, which is often overlooked in 2D calculations restricted to the plane only, the lowering in is so large (0.7Β eV) that the La-5 crosses the Ni-3 band and becomes the lowest lying conduction band at M.
+Beside this important qualitative modification of the Fermi surface, correlation corrections induce an even more evident and quantitative change in the bandwidth of the 5 band contributed by the two La spacer atoms only (i.e., with no contribution by the La inner atom lying in between the NiO planes, see Fig.
+13).
+This La-5 band, which in DFT-PBE is well above the Ni-3 one, is lowered by corrections, as it can be seen at in Fig.
+1d.
+At the M point, which is often overlooked in 2D calculations restricted to the plane only, the lowering in is so large (0.7 eV) that the La-5 crosses the Ni-3 band and becomes the lowest lying conduction band at M.
 Due to this large correction, the bottom of conduction (BOC) of the La-5 band at M is situated just only 30 meV above the top of valence (TOV) of the Ni-3 band at P (in the following, we will refer to this TOV-BOC distance of 30 meV as the indirect gap).
@@
 Note that a possible role of rare-earth self-doping was suggested [45] as responsible for the suppression of charge order, thereby possibly favoring superconductivity.
-We can expect that any tiny perturbation of the system, e.g.Β doping, vacancies or defects, can drive the system to a qualitatively different electronic structure and Fermi surface, and therefore affect the superconductivity.
-Note that even computationally, these low-energy features are highly sensitive to the uncertainty of the integration method to calculate the Fermi level, which we discuss in further detail in App.Β C.
+We can expect that any tiny perturbation of the system, e.g.
+doping, vacancies or defects, can drive the system to a qualitatively different electronic structure and Fermi surface, and therefore affect the superconductivity.
+Note that even computationally, these low-energy features are highly sensitive to the uncertainty of the integration method to calculate the Fermi level, which we discuss in further detail in App.
+D.
 We mention two recent works which pointed out the shift of the La-5 band due to correlations [15] and highlight the importance of this band on the screening which affects correlations on the Ni- bands [70].
 The latter work remarked in particular some coincidences between the lowering trend of the La-5 band with pressure and the shape of the superconducting region.
-In our calculation this band reaches the Fermi level at lower pressures and crosses the Fermi level already at 30 GPa at the M point, in comparison to 80 Gpa at the point in Verraes et al.Β [70].
+In our calculation this band reaches the Fermi level at lower pressures and crosses the Fermi level already at 30 GPa at the M point, in comparison to 80 Gpa at the point in Verraes et al.
+[70].
 III.1.3 Upper-lower Ni-3 energy gap
 A feature of the electronic structure that is considered of particular interest in relation to the emergence of superconductivity in La3Ni2O7 is the energy difference between the upper and lower Ni-3 bands.
-In the literature they are often labeled as bonding- and anti-bonding (as we explain in App.Β H, we find that this labeling is not appropriate).
+In the literature they are often labeled as bonding- and antibonding (as we explain in App.
+I, we find that this labeling is not appropriate).
 This energy difference is related to the Ni interlayer hopping.
-According to the scenario proposed by Sakakibara et al.Β [24], a superconductivity gap would open on these two Ni-3 bands even though they can be as far as 1Β eV away from the Fermi level, while the SC gap would be almost zero for the more Ni-3 character bands crossing the Fermi level.
+According to the scenario proposed by Sakakibara et al.
+[24], a superconductivity gap would open on these two Ni-3 bands even though they can be as far as 1 eV away from the Fermi level, while the SC gap would be almost zero for the more Ni-3 character bands crossing the Fermi level.
 In this scenario, a key parameter is precisely this Ni-3 interlayer hopping.
-As it can be seen in Fig.Β 1d, the two Ni-3 are practically unaffected by corrections, as their band width and separation remain essentially the same as in the DFT calculation.
-On this point our calculation is in disagreement both with the calculation of Christiansson et al.Β [45], which found a negative downward shift of the upper band of Β eV at the BZ corner, and with the calculation of You et al.Β [15], which in contrast found a positive upward shift of Β eV.
-Accordingly, our results leave practically unaffected the scenario proposed by Sakakibara et al.Β [24], while it would be boosted or disrupted in the two other cases.
+As it can be seen in Fig.
+1d, the two Ni-3 are practically unaffected by corrections, as their band width and separation remain essentially the same as in the DFT calculation.
+On this point our calculation is in disagreement both with the calculation of Christiansson et al.
+[45], which found a negative downward shift of the upper band of eV at the BZ corner, and with the calculation of You et al.
+[15], which in contrast found a positive upward shift of eV.
+Accordingly, our results leave practically unaffected the scenario proposed by Sakakibara et al.
+[24], while it would be boosted or disrupted in the two other cases.
 III.1.4 quasiparticle spectral weight and 1D character
 A calculation can also provide the quasiparticle spectral weight which represents a measure of the degree of correlation of the system.
-In Fig.Β 2a we show the interpolated bandplot of La3Ni2O7 at 29.5Β GPa, with the dots corresponding to the energies sampled on the -mesh.
+In Fig.
+2a we show the interpolated bandplot of La3Ni2O7 at 29.5 GPa, with the dots corresponding to the energies sampled on the -mesh.
 The relative size of the dots indicates the corresponding value of .
-In Fig.Β 2b we show the general orbital character of the bands (for more detailed projections, see Fig.Β 12).
-These plots are complemented with Fig.Β 3 which illustrates the behavior of as a function of the energy for the N -point of the Brillouin zone, a point we have chosen for the clear picture it gives.
-We note that La-5 and 4 bands, which at the level are difficult to distinguish, present the maximum QP weight culminating at .
+In Fig.
+2b we show the general orbital character of the bands (for more detailed projections, see Fig.
+15).
+These plots are complemented with Fig.
+3 which illustrates the behavior of as a function of the energy.
+We note that La-5 bands, followed by the 4, present the maximum QP weight culminating at .
 This is already quite remarkable, in particular for the most localized flat bands.
-Some of the La bands, in particular the previously discussed La-5 which almost achieved the Fermi level, bring these high spectral weights lower in energy than the main La manifold located around 5Β eV.
-For La-5 states above 5 eV, the QP weight is slightly decreasing (see Fig.Β 3).
-On the other side of the Fermi level, we also observe for the O-2 manifold below Β eV a reduced with increasing distance from with values ranging from 0.75 to 0.70.
+Some of the La bands, in particular the previously discussed La-5 which almost achieved the Fermi level, bring these high spectral weights lower in energy than the main La manifold located around 5 eV.
+For La-5 states above 5 eV, the QP weight is slightly decreasing.
+On the other side of the Fermi level, we also observe for the O-2 manifold below eV a reduced with increasing distance from with values ranging from 0.75 to 0.70.
 This is surprisingly lower than for La-5 and 4 states.
 However, the most remarkable fact is that the Ni-3 bands, with no difference between and , present the lowest weight of all quasiparticle states.
-For these states we observe a non-negligible drop of 0.15 when compared in particular with the La-5 which at M is situated at the same Fermi level as the Ni-3.
-This fact points to a larger weight of the non-coherent part of the spectrum for the Ni-3 states.
-We remind that in a Fermi liquid [59] the general trend is a reduction of the quasiparticle weight when going from the Fermi level to lower energies.
-Indeed, when going to the lowest energy quasiparticle states, the non-coherent part of spectra due to, e.g.Β plasmon losses, should increase, and correspondingly decreases.
-The same happens in the other higher energies direction, although not monotonically at the beginning.
-From this, we can conclude that already with the correlations introduced by a non-self-consistent approximation, La3Ni2O7 show a Fermi liquid trend for La-5, La-4 and O-2 quasiparticle states, but not for the Ni-3 ones around the Fermi level.
-The behaviour of the latter shows a departure from the Fermi liquid trend, but the minimal value of is still large enough not to configure a full breakdown of the Fermi liquid towards a strongly-correlated picture.
-As shown in Figs.Β 1a and b, the Fermi surface of La3Ni2O7 exhibits negligible dispersion along the direction, highlighting the pronounced 2D character of its electronic structure.
+For these states we observe a nonnegligible drop of up to 0.15 when compared in particular with the La-5 which at M is situated at the same Fermi level as the Ni-3.
+This fact points to a larger weight of the noncoherent part of the spectrum for the Ni-3 states.
+For comparison, in Fig.
+3 we also report the of the jellium model as in the Hedin calculation at and 2 [59], as well as at the La3Ni2O7 nominal average density of at 29.5 GPa ( with a.u., the unit cell volume, and the number of correlated electrons entering into the self-energy from the chosen pseudopotentials).
+The general trend for a Fermi liquid is a reduction of the quasiparticle weight when going from the Fermi level to lower energies.
+Indeed, when going to the lowest energy quasiparticle states, the noncoherent part of spectra due to, e.g.
+plasmon losses, should increase, and correspondingly decreases.
+The same happens in the other higher energies direction, although not monotonically at the beginning (for and 1.29, this decrease of occurs outside the plot range).
+From this, we can conclude that already with the correlations introduced by a non self-consistent approximation, La3Ni2O7 show a Fermi liquid trend for all quasiparticle states.
+We can associate the O-2 and even the La-5 and La-4 quasiparticle weights to the curve, that is to a higher level of correlation with respect to the system nominal density.
+On their side, the Ni-3 states show a departure from this Fermi liquid trend and so a even higher degree of correlation, although the minimal value of is still large enough not to configure a full breakdown of the Fermi liquid towards a strongly correlated picture.
+Of course, this is the picture emerging from a calculation which could be confuted by future experimental measures of the .
+Should a Fermi liquid breakdown be measured, we will have a failure of the approximation pointing to the need to go beyond it.
+As shown in Figs.
+1a and b, the Fermi surface of La3Ni2O7 exhibits negligible dispersion along the direction, highlighting the pronounced 2D character of its electronic structure.
 We note that, while this situation is similar to what is found in most of the cuprates, it differs from the infinite-layer nickelates case that presents large dispersion which introduces 3D features in their Fermi surface [71, 72].
@@
 These effectively 1D Fermi surfaces is an important difference with respect to cuprates.
-In particular, the emergence of or spin-density-wave ordersβas opposed to the antiferromagnetic order characteristic of the cupratesβcan be related to their nesting properties [34].
-In the cuprates, ARPES experiments have revealed the emergence of so-called βFermi arcsβ.
+In particular, the emergence of or spin-density-wave orders—as opposed to the antiferromagnetic order characteristic of the cuprates—can be related to their nesting properties [34].
+In the cuprates, ARPES experiments have revealed the emergence of so-called “Fermi arcs”.
 Specifically, the quasiparticle spectral weight is found to be locally suppressed at points along the Fermi surface that would coincide with the 1D Fermi surface segments.
@@
 III.2 Evolution with pressure
-In order to gain further insight on the key aspects that might be relevant for superconductivity in La3Ni2O7, we also analyze the evolution of its electronic structure with pressure.
-We remind that La3Ni2O7 enters into a superconducting phase at an experimentally measured pressure of 14Β GPa without doping [1].
-To explore the largest range, we have included in our study a fictitious zero pressure phase of La3Ni2O7 enforced in the tetragonal I4/mmm crystal structure, which in reality at 0Β GPa should be unstable towards the orthorhombic Amam phase with tilted octahedra.
+To gain further insight on the key aspects that might be relevant for superconductivity in La3Ni2O7, we also analyze the evolution of its electronic structure with pressure.
+We remind that La3Ni2O7 enters into a superconducting phase at an experimentally measured pressure of 14 GPa without doping [1].
+To explore the largest range, we have included in our study a fictitious zero pressure phase of La3Ni2O7 enforced in the tetragonal I4/mmm crystal structure, which in reality at 0 GPa should be unstable towards the orthorhombic Amam phase with tilted octahedra.
 Of course, at the present knowledge we cannot exclude that the structural transition occurs at the same critical pressure as the transition to superconductivity, so that it is an important ingredient to explain the superconductivity in this system.
-In this case, our calculation of the tetragonal phase at 0Β GPa would be probably inadequate to study the normal-superconducting phase transition.
+In this case, our calculation of the tetragonal phase at 0 GPa would be probably inadequate to study the normal-superconducting phase transition.
 But beyond the fact that a calculation is less cumbersome in systems with more symmetries, our choice is motivated by the fact that the comparison of electronic structures is difficult between different crystal structures.
 Topological variations in the Fermi surface might trigger the superconducting phase transition (and maybe also the structural one): the approach proposed in this section focus on them.
-In Fig.Β 4 we compare the electronic structure of La3Ni2O7 at ambient pressure and at 29.5Β GPa (in the same figure we also report the band plot of a biaxial strained La3Ni2O7 which will be discussed in the next section).
-The most remarkable difference between the two pressures is in the La-5 band at the point M which, at ambient pressure, lies eV above the upper Ni-3 band in a very similar way to what already seen but at the DFT-PBE level in the 29.5 GPa case (see Fig.Β 1d), though in the latter the effect is much larger (0.5 eV).
+In Fig.
+4 we compare the electronic structure of La3Ni2O7 at ambient pressure and at 29.5 GPa (in the same figure we also report the band plot of a biaxial strained La3Ni2O7 which will be discussed in the next section).
+The most remarkable difference between the two pressures is in the La-5 band at the point M which, at ambient pressure, lies eV above the upper Ni-3 band in a very similar way to what already seen but at the DFT-PBE level in the 29.5 GPa case (see Fig.
+1d), though in the latter the effect is much larger (0.5 eV).
 That is, correlations and pressure act in the same direction towards the emergence of the La-5 band at the lowest energies.
-For a more complete discussion on this point, see App.Β B.
+For a more complete discussion on this point, see App.
+B.
 At ambient pressure the La-5 band is not present in the low energy region, and therefore the delicate balance between the two incipient self-doping bands is lost.
 The electronic structure of La3Ni2O7 at ambient pressure has been measured by an ARPES experiment [56] which we can now compare with our calculation at zero GPa.
-The hole pocket, which is found in the DFT PBE Fermi surface as due to the lower Ni-3 band, is absent in the ARPES experiment, exactly like in (Fig.Β 1b and Fig.Β 16 in the Appendix).
-ARPES spectra situate the top of this band at the BZ corners and 50Β meV below the Fermi level.
-Our calculation at zero GPa (see Fig.Β 4), found the top of this band 50Β meV below the Fermi level at the X -point (corresponding to in Yang et al.Β [56], see Fig.Β 1c for -points correspondence).
-In contrast to the La-5 band, this lower Ni-3 band is almost unaffected by hydrostatic pressure: in the results at 29.5Β GPa it just only shifts down Β meV, which validates the comparison of the latter case with the ARPES, even if not performed at the same pressure.
-In Fig.Β 4b, we report a graph presenting the top-of-valence (TOV) and bottom-of-conduction (BOC) positions of the most relevant bands for different hydrostatic pressures.
-We report extra calculations done at 14 and at 40Β GPa, covering the range of superconductivity as measured experimentally in Sun et al.Β [1].
-The indirect gap is at its minimum around 29.5Β GPa, and the La-5 band is clearly self-doping beyond.
-From the same figure we can also remark that the same La-5 band crosses the Ni-3 upper band precisely at 14Β GPa, which is precisely the onset of the superconducting region measured by Sun et al.Β [1].
+The hole pocket, which is found in the DFT PBE Fermi surface as due to the lower Ni-3 band, is absent in the ARPES experiment, exactly like in (Fig.
+1b and Fig.
+17 in the Appendix).
+ARPES spectra situate the top of this band at the BZ corners and 50 meV below the Fermi level.
+Our calculation at zero GPa (see Fig.
+4), found the top of this band 50 meV below the Fermi level at the X -point (corresponding to in Yang et al.
+[56], see Fig.
+1c for -points correspondence).
+In contrast to the La-5 band, this lower Ni-3 band is almost unaffected by hydrostatic pressure: in the results at 29.5 GPa it just only shifts down meV, which validates the comparison of the latter case with the ARPES, even if not performed at the same pressure.
+In Fig.
+4b, we report a graph presenting the top-of-valence (TOV) and bottom-of-conduction (BOC) positions of the most relevant bands for different hydrostatic pressures.
+We report extra calculations done at 14 and at 40 GPa, covering the range of superconductivity as measured experimentally in Sun et al.
+[1].
+The indirect gap is at its minimum around 29.5 GPa, and the La-5 band is clearly self-doping beyond.
+From the same figure we can also remark that the same La-5 band crosses the Ni-3 upper band precisely at 14 GPa, which is precisely the onset of the superconducting region measured by Sun et al.
+[1].
 This might be just only a coincidence or have deeper implications.
+(Note that the position of this crossing and the overall picture does not change by performing a calculation in the ambient and low pressure Amam structure which will place the position of this band at even larger energy with respect to its I4/mmm position shown in Fig.
+4 at 0 GPa, see App.
+C).
 In support to the latter, we cite a very recent work [70] which also found a correlation between the position of the La-5 band with respect to the Fermi level, and the shape of the superconducting region.
-Their calculation is at the DFT level, so that they missed the earlier coincidence in pressure of the La-5 with the Ni-3 band precisely at the onset of superconductivity at 14Β GPa, but the relevance of the La-5 for superconductivity has been remarked.
+Their calculation is at the DFT level, so that they missed the earlier coincidence in pressure of the La-5 with the Ni-3 band precisely at the onset of superconductivity at 14 GPa, but the relevance of the La-5 for superconductivity has been remarked.
 We can more carefully conclude that both correlations and pressure are fundamental to describe the electronic structure of La3Ni2O7 in its superconducting phase.
@@
 III.3 La3Ni2O7 under epitaxial strain
-Recent works [5, 6] have reported superconductivity at Β K in La3Ni2O7 thin films under compressive strain at ambient pressure.
-Here we present the electronic structure of a bulk La3Ni2O7 with a 1.8% strained Β Γ Β and a relaxed , which we use to simulate the experimental thin film.
-The corresponding PBE and band plots, PDOS and DOS are shown in Fig.Β 5.
-In comparison to the previously discussed case at ambient and 29.5Β GPa (see also Fig.Β 4), the electronic structure retains most of its features.
+Recent works [5, 6] have reported superconductivity at K in La3Ni2O7 thin films under compressive strain at ambient pressure.
+Here we present the electronic structure of a bulk La3Ni2O7 with a 1.8% strained Å and a relaxed , which we use to simulate the experimental thin film.
+The corresponding PBE and band plots, PDOS and DOS are shown in Fig.
+5.
+In comparison to the previously discussed case at ambient and 29.5 GPa (see also Fig.
+4), the electronic structure retains most of its features.
 The Fermi level crossing is practically the same in the three different cases.
-In particular, the cuprate-like sheet has exactly the same shape (refer to Fig.Β 1).
+In particular, the cuprate-like sheet has exactly the same shape (refer to Fig.
+1).
 The lower Ni-3 band is significantly further from in the thin film case ( eV).
@@
 First, the hole pocket is already absent at the DFT level, that is the lower Ni-3 band does not cross any longer, an effect that is further enhanced by .
-Second, the La-5 band at M, which also undergoes a large negative correction of almost 1Β eV, now stays above the Fermi level by eV.
+Second, the La-5 band at M, which also undergoes a large negative correction of almost 1 eV, now stays above the Fermi level by eV.
 In fact, it only reaches the upper Ni-3 band without crossing it.
-We can see that, with respect to the band positions at M, the epitaxial strained configuration is intermediate between 29.5Β GPa and ambient pressure.
-Although our biaxial strain simulation implements a non-hydrostatic realization, we can compare it to the hydrostatic pressure case where the La-5 BOC is degenerate with Ni-3 BOC at M.
-Following this analogy, from Fig.Β 4b, where we have also reported the band positions for the thin film, we can tentatively attribute our epitaxial strain simulation as equivalent to the 14Β GPa hydrostatic pressure case, that is precisely at the superconducting onset.
+We can see that, with respect to the band positions at M, the epitaxial strained configuration is intermediate between 29.5 GPa and ambient pressure.
+Although our biaxial strain simulation implements a nonhydrostatic realization, we can compare it to the hydrostatic pressure case where the La-5 BOC is degenerate with Ni-3 BOC at M.
+Following this analogy, from Fig.
+4b, where we have also reported the band positions for the thin film, we can tentatively attribute our epitaxial strain simulation as equivalent to the 14 GPa hydrostatic pressure case, that is precisely at the superconducting onset.
 This could be an indication that the relative position and overlap of BOC at M, might play a role as a triggering mechanism of superconductivity.
@@
 It is claimed that, in addition to the and sheets, the pocket is also present at the Fermi level.
-In another more recent ARPES experiment on a La2PrNi2O7 thin film [73], the pocket is absent from the Fermi level and the Ni-3 band is found 70Β meV below the Fermi energy, which compares more favorably with our result of 200Β meV.
+In another more recent ARPES experiment on a La2PrNi2O7 thin film [73], the pocket is absent from the Fermi level and the Ni-3 band is found 70 meV below the Fermi energy, which compares more favorably with our result of 200 meV.
 In any case, the comparison with our electronic structure is difficult because they considered a Pr doped system.
@@
 Our results should then be considered as an ab initio prediction of the electronic structure which could be measured for a La3Ni2O7 thin film of enough large thickness.
-Or, possibly, for a non-hydrostatic planar strain applied on a bulk sample that can be realized experimentally by applying a tensile strain to the axis, which will induce a biaxial compressive strain of the plane by Poisson law.
-We can conclude that our calculation on the epitaxial strained case presents an electronic structure very similar to the bulk at 14Β GPa pressure.
+Or, possibly, for a nonhydrostatic planar strain applied on a bulk sample that can be realized experimentally by applying a tensile strain to the axis, which will induce a biaxial compressive strain of the plane by Poisson law.
+We can conclude that our calculation on the epitaxial strained case presents an electronic structure very similar to the bulk at 14 GPa pressure.
 Correlation effects are very important on both the lower Ni-3 band, causing the disappearance of the pocket, and on the increased relevance of the La-5 band at low energy.
@@
 Therefore, the Fermi surface is composed only of the cuprate-shape sheet plus the nickelate-specific cylinder , both showing an effective 1D character, and is in good agreement with ARPES experiments [56] which did not find any pocket.
-Additionally, we observe a non-negligible drop in the quasiparticle spectral weight on the Ni-3 states, both and , so to configure a little departure from a Fermi liquid behavior.
+Additionally, we observe a nonnegligible drop in the quasiparticle spectral weight on the Ni-3 states, both and , so to configure a little departure from a Fermi liquid behavior.
 Finally, we have shown that not only correlations, but also pressure or biaxial strain play a crucial role in pulling down the La-5 band at and M towards the Fermi level, which may induce self-doping.
@@
 V Acknowledgments
-We thank Markus Holzmann for useful discussions. We acknowledge financing from the LABEX LANEF project NICOS. Computing time has been provided by French GENCI and Grenoble CIGRID.
-Appendix A Crystal structures used in the calculations
-TAB.Β 1 reports the crystal structures in the conventional cell used for the calculation at 29.5Β GPa and for the in-plane strained system.
-In these two cases the parameters were chosen with a trade off between the values from the literature and obtained via a structural optimization.
-TAB.Β 2 reports the crystal structures used for the extra pressure calculations, all obtained via structural optimizations.
-| 29.5 GPa [1, 24] | a=3.7148 | c=19.7340 | ||
-| Ni | 0 | 0 | 0.096 | |
-| La1 | 0 | 0 | 0.321 | |
-| La2 | 0 | 0 | 0.5 | |
-| O1 | 0 | 0.5 | 0.095 | |
-| O2 | 0 | 0 | 0.204 | |
-| O3 | 0 | 0 | 0 | |
-| epitaxial strained (thin film at ambient pressure) [5] | a=3.77 | c=20.6927 | ||
-| Ni | 0 | 0 | 0.095 | |
-| La1 | 0 | 0 | 0.321 | |
-| La2 | 0 | 0 | 0.5 | |
-| O1 | 0 | 0.5 | 0.096 | |
-| O2 | 0 | 0 | 0.205 | |
-| O3 | 0 | 0 | 0 | |
+We thank Markus Holzmann for useful discussions.
+We acknowledge financing from the LABEX LANEF project NICOS.
+Computing time has been provided by French GENCI Grant No.
+2022-AD01091394 and Grenoble CIGRID.
+| 29.5 GPa [1, 24] | a=3.7148 | c=19.7340 | ||||
+| Ni | 0 | 0 | 0.096 | |||
+| La1 | 0 | 0 | 0.321 | |||
+| La2 | 0 | 0 | 0.5 | |||
+| O1 | 0 | 0.5 | 0.095 | |||
+| O2 | 0 | 0 | 0.204 | |||
+| O3 | 0 | 0 | 0 | |||
+a=3.77 | c=20.6927 | ||||
+| Ni | 0 | 0 | 0.095 | |||
+| La1 | 0 | 0 | 0.321 | |||
+| La2 | 0 | 0 | 0.5 | |||
+| O1 | 0 | 0.5 | 0.096 | |||
+| O2 | 0 | 0 | 0.205 | |||
+| O3 | 0 | 0 | 0 | |||
 | 0 GPa | a=3.8408 | c=20.2929 | ||
@@
 | O3 | 0 | 0 | 0 | |
+Appendix A Crystal structures used in the calculations
+TAB. 1 reports the crystal structures in the conventional cell used for the calculation at 29.5 GPa and for the in-plane strained system.
+In these two cases the parameters were chosen with a trade off between the values from the literature and obtained via a structural optimization.
+TAB. 2 reports the crystal structures used for the extra pressure calculations, all obtained via structural optimizations.
 Appendix B Sensitiveness of the electronic structure to internal atomic positions
@@
 The latter is then further modified by the variation of the lattice parameters under pressure which is fundamental to trigger superconductivity in La3Ni2O7.
-In our first attempt to calculate the DFT and the electronic structures of the bilayer at 29.5 GPa, we used the experimental atomic positions determined by X-ray diffraction (XRD) given in the extended data table I in Sun et al.Β [1] for the Fmmm phase and adapted to the as also done in Sakakibara et al.Β [24]
-The resulting DFT and Fermi surfaces (Fig.Β 6) show the appearance of a cylindrical Fermi sheet centered on which is not observed in the ARPES experiment of Yang et al.Β [56] This fake Fermi sheet originates from the crossing at the Fermi level of the La-5 character band, as one can see in Fig.Β 7a for PBE and in Fig.Β 7b for .
+In our first attempt to calculate the DFT and the electronic structures of the bilayer at 29.5 GPa, we used the experimental atomic positions determined by X-ray diffraction (XRD) given in the extended data table I in Sun et al.
+[1] for the Fmmm phase and adapted to the as also done in Sakakibara et al.
+[24]
+The resulting DFT and Fermi surfaces (Fig.
+6) show the appearance of a cylindrical Fermi sheet centered on which is not observed in the ARPES experiment of Yang et al.
+[56] This fake Fermi sheet originates from the crossing at the Fermi level of the La-5 character band, as one can see in Fig.
+7a for PBE and in Fig.
+7b for .
 It has to be considered as an artifact due to a non correct configuration of the internal atomic positions.
@@
 II.2 of the main text.
-The DFT bandplot of this relaxed structure is shown on Fig.Β 7c and has to be compared with Fig.Β 7a corresponding to the experimental structure.
+The DFT bandplot of this relaxed structure is shown on Fig.
+7c and has to be compared with Fig.
+7a corresponding to the experimental structure.
 In particular, we remark that the self-doping induced by the La-5 band is absent for the relaxed structure.
@@
 To this purpose we computed (at the DFT level only) the bandplot of a structure similar to the relaxed one, but with the La atoms in the spacer (La1 in Tab.
-1) artificially shifted away from the NiO2 plane by 0.14Β Γ , which corresponds to 0.7% of (see Fig.
+1) artificially shifted away from the NiO2 plane by 0.14 Å, which corresponds to 0.7% of (see Fig.
 8).
 The associated bandplot is shown on Fig.
-7d and reproduces well the self-doping behavior of the rare-earth band that is observed on the DFT bandplot of Fig.Β 7a, and to a lesser extent on the bandplot of Fig.Β 1d.
+7d and reproduces well the self-doping behavior of the rare-earth band that is observed on the DFT bandplot of Fig.
+7a, and to a lesser extent on the bandplot of Fig.
+1d.
 We could have rather moved the apical oxygens of the spacer toward the NiO2 planes, but this would have changed the shape of the octahedron and potentially the crystal field splitting.
@@
 We hope that future experiments, for example using other techniques like neutron diffraction measures, could be helpful in clarifying the internal atomic positions which are critical for the electronic structure.
-We anticipate that and we are working on a more complete study [74] about this effect and its correct description, both the internal atomic positions and the numerical methods employed being important.
-This can further help shading light on the aspects which could be relevant to superconductivity in RP nickelates.
-Appendix C Fermi energy calculation: standard bisection vs DOS integration methods
+Appendix C vs band plots
+In this Section we report the band plot for the Amam crystal structure relaxed at GPa, shown in its oS base-centered orthorhombic BZ (Fig.
+9 red lines).
+For comparison we also show the band plot for the 0 GPa relaxed system but constrained within the I4/mmm crystal structure and folded within the same oS BZ (blue lines) from its original tI body-centered tetragonal (BCT) BZ.
+Of course, the two BZ have different topology and even after the folding they do not overlap.
+High symmetry points of the tI may not correspond to any high symmetry point of the oS.
+And even in case of one-to-one correspondence between high symmetry points, they might be located at slightly different positions due to the different BZ shape determined by the to (and to ) ratios which of course differ.
+We remark some differences, especially in the conduction states which however are not observed by direct ARPES photoemission, due to the reduction of symmetry induced by the tilting of NiO6 octahedra and by the introduced orthorhombicity which are both at their maximum at 0 GPa.
+Nevertheless, the two plots are very similar at the Fermi level.
+In particular the presence of the hole pocket is confirmed also in the DFT PBE Amam structure: only folded from the X-P segment of the BCT Brillouin zone, into the (initial part of the) -Y segment of the oS BZ.
+The only noticeable difference is, in the Amam structure, a slightly larger dispersion of the band along the axis than in the I4/mmm where it is almost undispersive.
+The Amam band is either at the same level (in ) or shifted meV downward (in Y) compared to the I4/mmm one.
+We know that the general tendency of the corrections is to shift down this band by meV with respect to DFT PBE.
+Therefore the pocket should disappear also in a calculation of the Amam structure.
+Finally, outside the Fermi region we remark that the La-5 band is further shifted up in the Amam crystal structure with respect to the I4/mmm where, as discussed in the main text, it is found a little below 1 eV at the M point (corresponding to the Y point of the oS).
+Therefore this band is further outside the low-energy region for the combined effect of pressure and symmetry reduction, and enter into play at the pressure of 14 GPa, as illustrated in Fig.
+4b.
+This comparison indicates that the main conclusions of this work, and in particular about the topology of the Fermi surface and the disappearance of the hole pocket, should also hold for the case of the Amam crystal structure at ambient pressure as measured by ARPES.
+Appendix D Fermi energy calculation: standard bisection vs DOS integration methods
 The indirect gap of 30 meV between the two incipient bands found in the calculation done at 29.5 GPa is beyond the accuracy by which the Fermi level can be calculated.
-This is already a consequence of the finite smearing temperature used to calculated metals, of the order of Β Ha Β meV and up to Β Ha Β meV (the Abinit default is Β Ha, while it is half that in Quantum Espresso).
-We must also keep in mind that any change in the parameters used to generate the pseudo-potentials can affect DFT Kohn-Sham energies by even 100200Β meV.
-This is also the reason why the quasiparticle energy convergence is never pushed beyond 100Β meV.
+This is already a consequence of the finite smearing temperature used to calculated metals, of the order of Ha meV and up to Ha meV (the Abinit default is Ha, while it is half that in Quantum Espresso).
+We must also keep in mind that any change in the parameters used to generate the pseudopotentials can affect DFT Kohn-Sham energies by even 100200 meV.
+This is also the reason why the quasiparticle energy convergence is never pushed beyond 100 meV.
 A further indetermination of the Fermi energy arise in particular from the choice of the method used to calculate it.
-In this Section we present the comparison between two of them: the classical method using a bisection algorithm, as implemented in Abinit or Quantum Espresso; and the Wannier interpolation as implemented in Wannier90, followed by the integration of the total DOS done with our own Python code using the Simpson method.
+In this Section we present the comparison between two of them: the classical method using a bisection algorithm, as implemented in Abinit or Quantum Espresso; and the integration of the Wannier90 total DOS done with our own Python code using the Simpson method.
+The Wannier90 DOS is calculated from the PWF-interpolation bands, on a grid.
 In this Section we show how critical the calculation of the Fermi energy can be on the electronic structure and the Fermi surface of this system.
-For this reason, we take the most critical case of 29.5Β GPa and we check the two methods here above described (for short labeled the bisection and DOS methods), keeping as close as possible to their defaults when there is matter of choice.
-In Fig.Β 9 we show a zoom around the Fermi level of the bandplot, as in Fig.Β 1d but reporting the different Fermi energies calculated using a smearing temperature of Ha, with the corresponding Fermi surfaces below.
-First of all, one can remark the robustness of the and sheets with respect to a large variation of (120 meV), which is due to the more dispersive character of these bands compared with to two incipient ones.
-With the bisection method ( traced by a black line and set as the origin), a pocket of electrons () appears around the M point in the plane (see BZ scheme on Fig.Β 1c), due to the La-5 band crossing.
+For this reason, we take the most critical case of 29.5 GPa and we check the two methods here above described (for short labeled the bisection and DOS methods), keeping as close as possible to their defaults when there is matter of choice.
+In Fig.
+10 we show a zoom around the Fermi level of the bandplot calculated using a smearing temperature of Ha, as in Fig.
+1d, but reporting the different Fermi energies calculated and the associated Fermi surfaces below.
+First of all, one can remark the robustness of the and sheets with respect to a large variation of (100 meV), which is due to the more dispersive character of the corresponding bands compared with to two incipient ones.
+With the bisection method ( traced by a black line and set as the origin), a pocket of electrons () appears around the M point in the plane (see BZ scheme on Fig.
+1c), due to the La-5 band crossing.
 On the other hand, with the DOS method and using the least arbitrary Wannier functions set among all possible sets, the PWF one, the Fermi energy is 87 meV (purple line) lower.
-In this case, there is no pocket of electrons at M and, in contrast, a pocket of holes is opened in the XP direction in the calculation, and the Fermi surface becomes again very similar to the DFT-PBE one, i.e.Β with the reappearance of the pocket in the BZ corners.
+In this case, there is no pocket of electrons at M and, in contrast, a pocket of holes is opened in the XP direction in the calculation, and the Fermi surface becomes again very similar to the DFT-PBE one, i.e.
+with the reappearance of the pocket in the BZ corners.
 As it can be seen, different methods provide qualitatively different Fermi surfaces leading to different low energy physics.
-In order to improve our estimate on the Fermi energy, we have performed a full DFT+ calculation reducing the smearing temperature to Ha Β meV.
-The corresponding quasiparticle energies are indicated by pink stars and the associated by the green dashed line in Fig.Β 9.
+In order to improve our estimate on the Fermi energy, we have performed a full DFT+ calculation reducing the smearing temperature to Ha meV.
+The corresponding quasiparticle energies are indicated by green stars and the associated by the green dashed line in Fig.
+10.
 As we can see, now the Fermi level situated more or less in the middle of the indirect band gap, at meV compare to the calculated with Ha.
 As a consequence both the electron pocket and the hole pocket are suppressed.
-And this is the Fermi level we have used to plot the Fermi surface in Fig.Β 1b.
+And this is the Fermi level we have used to plot the Fermi surface in Fig.
+1b.
 In practice, in a system with such a delicate balance, any defect, vacancy (La3Ni2O7 is usually oxygen deficient) or impurity, introduces a tiny perturbation in the Fermi energy which will open the electron or the hole pocket.
-In any case, the Fermi surface showed on Fig.Β 1b corresponds to a theoretical ideal situation.
-Therefore, we expect that a real experiment will probably not measure the ideal situation.
-Therefore we leave a margin of error of 50Β meV on the Fermi level to keep into account both experimental and theoretical uncertainties.
-Notice that 29.5Β GPa is a really critical pressure for the La3Ni2O7 electronic structure.
-Indeed, at ambient pressure the La-5 band is almost 0.5Β eV above its position at 29.5Β GPa (see Fig.Β 4b), and the bisection method set the Fermi energy well below the bottom of this band, with no possibility of an electron pocket opened at M.
-Consequently, the Fermi surface would be exactly that in Fig.Β 1b, with the stable and sheets and no or pockets at all, from ambient pressure on, and without any relevant modification up to 29.5Β GPa.
-Appendix D Orbital character projection: pure atomic orbitals vs Wannier functions
+In any case, the Fermi surface showed on Fig.
+1b corresponds to a theoretical ideal situation.
+Therefore, we expect that a real experiment will probably not measure the ideal situation, and we leave a margin of error of 100 meV on the Fermi level to keep into account both experimental and theoretical uncertainties.
+Notice that 29.5 GPa is a really critical pressure for the La3Ni2O7 electronic structure.
+Indeed, at ambient pressure the La-5 band is almost 0.5 eV above its position at 29.5 GPa (see Fig.
+4b), and the bisection method set the Fermi energy well below the bottom of this band, with no possibility of an electron pocket opened at M.
+Consequently, the Fermi surface would be exactly that in Fig.
+1b, with the stable and sheets and no or pockets at all, from ambient pressure on, and without any relevant modification up to 29.5 GPa.
+Appendix E Orbital character projection: pure atomic orbitals vs Wannier functions
 Fig.
-10 shows three DFT PBE band character plots of La3Ni2O7 at 29.5 GPa: the first one using projections on (pseudo) atomic orbitals calculated with Quantum Espresso, which we will consider as the reference; and the two others with Wannier90 but using different WF sets, namely projected Wannier functions (PWF) and maximally-localized Wannier functions (MLWF).
+11 shows three DFT PBE band character plots of La3Ni2O7 at 29.5 GPa: the first one using projections on (pseudo) atomic orbitals calculated with Quantum Espresso, which we will consider as the reference; and the two others with Wannier90 but using different WF sets, namely projected Wannier functions (PWF) and maximally localized Wannier functions (MLWF).
 PWF are the Wannier90 first guess and correspond to WF with imposed well defined orbital character (in the input file), though not purely hydrogenoid atomic orbitals [74, 75] (see below).
@@
 The Quantum Espresso and the PWF band plot characters are very consistent between themselves, whereas the MLWF shows important discrepancies for the La- orbital character which lose weight on the band of interest and is more hybridized with the 4 manifold.
-It must be kept into account that the projection onto MLWF, beyond to strongly depend on the chosen set of MLWF (e.g.Β localized on atoms, on bonds, or on anti-bonding positions), is different from the projection onto atomic wave functions.
+It must be kept into account that the projection onto MLWF, beyond to strongly depend on the chosen set of MLWF (e.g.
+localized on atoms, on bonds, or on antibonding positions), is different from the projection onto atomic wave functions.
 But there is large arbitrariness also for atomic wave functions.
-For example whether projecting onto all-electron wave functions (as in all-electron codes), or pseudo-atomic wave functions (e.g.Β Quantum Espresso and Abinit, with some minor differences between them), or spherical harmonics (e.g.Β VASP).
+For example whether projecting onto all-electron wave functions (as in all-electron codes), or pseudoatomic wave functions (e.g.
+Quantum Espresso and Abinit, with some minor differences between them), or spherical harmonics (e.g.
+VASP).
 In conclusion, projections can be a convenient label to identify bands, establishing a convention.
@@
 Note that this orbital character comparison between Quantum Espresso and Wannier90 has been done for the PDOS as well, providing the same conclusion.
-Appendix E Additional orbital character projections
-On Fig.Β 12 we provide more precise band character plots than on Fig.Β 2b, giving details on the specific orbitals involved in the Bloch states.
-We distinguish the two spacer La atoms from the inner one lying in between the two NiO2 planes; as well as the planar O atoms lying in the NiO2 planes, and the apical O atoms of the spacer (ap.Β sp.) or of the inner layer (ap.Β inner).
-See Fig.Β 13 which shows the atoms labeling on the Vesta structure.
+One can remark on Fig.
+12 that PWF do not correspond to real atomic orbitals, showing extra lobes and adaptation to neighboring atoms.
+This is due to the projection onto Bloch states and to the Löwdin orthogonalization procedure [75] as described in Souza et al.
+[74]
+Appendix F Additional orbital character projections
+On Fig.
+15 we provide more precise band character plots than on Fig.
+2b, giving details on the specific orbitals involved in the Bloch states.
+We distinguish the two spacer La atoms from the inner one lying in between the two NiO2 planes; as well as the planar O atoms lying in the NiO2 planes, and the apical O atoms of the spacer (ap.
+sp.) or of the inner layer (ap.
+inner).
+See Fig.
+13 which shows the atoms labeling on the Vesta structure.
 For O-2, we also distinguish and orbitals.
-We remark that there is a non-negligible La-4 orbital character on the lowered La-5 band which almost reaches the Fermi level.
+We remark that there is a nonnegligible La-4 orbital character on the lowered La-5 band which almost reaches the Fermi level.
 We also remark that in-plane O-2 orbitals have a strong contribution on what we label Ni-3 bands, especially above the Fermi level.
-Appendix F Convergence in the number of bands
+Appendix G Convergence in the number of bands
 In a calculation, the number of bands taken into account in the calculation of the self-energy is critical.
 A convergence study is therefore needed.
-Moreover, this study is motivated by the discrepancy on the Ni-3 upper and lower bands which we have found between our results and the results of Christiansson et al.Β [45] from one side, and of You et al.Β [15] from another, especially regarding the shift the latter obtained on the upper Ni-3 band from DFT to .
+Moreover, this study is motivated by the discrepancy on the Ni-3 upper and lower bands which we have found between our results and the results of Christiansson et al.
+[45] from one side, and of You et al.
+[15] from another, especially regarding the shift the latter obtained on the upper Ni-3 band from DFT to .
 In our original calculation, we used 225 bands.
 For this test we rise up to 300 for the calculation of energies of three bands around the Fermi level.
-The energies obtained are plotted on top of the original calculation in Fig.Β 14.
-As we can see, the original calculation is already enough converged, the maximal difference for the set of checked bands being of Β meV, validating the electronic structure presented in the main text.
-Appendix G Fermi surfaces of the substrate constrained and ambient pressure cases
-Appendix H Bloch state plots
+The energies obtained are plotted on top of the original calculation in Fig.
+14.
+As we can see, the original calculation is already enough converged, the maximal difference for the set of checked bands being of meV, validating the electronic structure presented in the main text.
+Appendix H Fermi surfaces of the substrate constrained and ambient pressure cases
+Appendix I Bloch state plots
 In this section we show some selected Bloch states that are interesting to visualize in real space.
 Notice that these Bloch states are calculated in DFT PBE, as our one-shot first order perturbation theory calculation does not recalculate the wavefunctions, but only corrects the energies.
-The lower and upper Ni-3 bands are sometimes respectively labeled as bonding and anti-bonding, according to a chemical convention.
-For these two states we plot in Fig.Β 17 the iso-surfaces of the Bloch wave functions at the X -point which lies in the corner of the BZ.
-As we can see, these names are not appropriated, the upper state being more βbondingβ than the lower one, with more weight on orbitals of inner oxygen atoms lying in between NiO2 planes.
+The lower and upper Ni-3 bands are sometimes respectively labeled as bonding and antibonding, according to a chemical convention.
+For these two states we plot in Fig.
+18 the iso-surfaces of the Bloch wave functions at the X -point which lies in the corner of the BZ.
+As we can see, these names are not appropriated, the upper state being more “bonding” than the lower one, with more weight on orbitals of inner oxygen atoms lying in between NiO2 planes.
 Another Bloch state worth to be shown is the one corresponding to the La-5 band, which is the most sensitive to pressure, and in particular at the M point where it reaches its minimum.
-In Fig.Β 18 we see that these electrons are delocalized in the spacer, which explain the highly dispersive shape of the corresponding band.
-From 0 to 29.5Β GPa, this state is not occupied.
-It starts to be occupied beyond 29.5Β GPa, inducing self-doping of the NiO2 planes.
+In Fig.
+19 we see that these electrons are delocalized in the spacer, which explain the highly dispersive shape of the corresponding band.
+From 0 to 29.5 GPa, this state is not occupied.
+It starts to be occupied beyond 29.5 GPa, inducing self-doping of the NiO2 planes.